### v1

# Evolving Collective Cognition in Human–Agent Hybrid Societies: How Agents Form Stances and Boundaries

Abstract
Large language models have been widely used to simulate credible human social behaviors. However, it remains unclear whether these models can demonstrate stable capacities for stance formation and identity negotiation in complex interactions, as well as how they respond to human interventions. We propose a computational multi-agent society experiment framework that integrates generative agent-based modeling with virtual ethnographic methods to investigate how group stance differentiation and social boundary formation emerge in human–Agent hybrid societies. Across three studies, we find that agents exhibit endogenous stances, independent of their preset identities, and display distinct tonal preferences and response patterns to different discourse strategies. Furthermore, through language interaction, agents actively dismantle existing identity-based power structures and reconstruct self-organized community boundaries based on these stances. Our findings suggest that preset identities do not rigidly determine the agents’ social structures. For human researchers to effectively intervene in collective cognition, attention must be paid to the endogenous mechanisms and interactional dynamics within the agents’ language networks. These insights provide a theoretical foundation for using generative AI in modeling group social dynamics and studying human–agent collaboration.
In recent years, large language models (LLMs) have made continuous advances in their ability to generate, understand, and respond to natural language, fostering the emergence of a new kind of human–machine sociology. In this emerging context, humans are no longer merely operators of models but co-inhabitants of interactive systems alongside agents capable of social perception and expression [1, 2, 3]. LLM-based are increasingly exhibiting the characteristics of “actors”[4, 5]. The simulation of social behavior has also undergone a fundamental shift, from encoding static rules and actions to the generative exploration of emergent social behavior itself [6]. This transformation gives rise to human–agent Hybrid Societies.
Within such systems, individual cognition is gradually formed through mechanisms such as interaction, imitation, response, and revision. These processes unfold in ongoing language exchanges and social engagement, shaped both by the values of human participants and by the propagation of language among agents [1]. This phenomenon corresponds to what is called group cognition: a dynamic system in which cognition is embedded within community relations, arising through discursive exchanges among group members and irreducible to preset stances held by individuals [7].
From this perspective, agents are not merely tools for task execution but become discursive participants in a shared communicative space. They are capable of forming attitudes, negotiating boundaries, and engaging with institutional structures through interactions with both humans and other agents. While existing studies often focus on the capabilities and controllability of agents [8], they have paid less attention to the generative roles agents may play in constructing social cognition. A key tension arises here: behavioral controllability often depends heavily on preset identities and stances assigned to agents. Many current studies of artificial sociality attempt to simulate society by assigning agents predefined roles and power relations [9, 8]. However, the theory of group cognition suggests that phenomena such as stance differentiation, group affiliation, and boundary negotiation emerge dynamically from interactions among diverse actors rather than from fixed designs [10].
This gap raises a fundamental question: when agents are embedded within social interactions, can they still form attitudes, express stances, and participate in boundary formation through language practices as their preset identities would suggest? Furthermore, can human researchers shape the evolution of collective cognition not through predefinition, but through embodied interventions and co-construction of shared meaning spaces? By building an integrated theoretical and experimental framework, we aim to better understand the formation mechanism of group cognition in a human-agent hybrid society. This understanding points toward a new kind of human–agent sociology [11].
1 The Current Project
Our research focus on the dynamic emergence of stance within agent collectives. Since research on human discourse politics often finds that statistical surveys alone are insufficient to capture the micro-level influence dynamics and the emergence of social cognition within discursive practices [12, 13], and existing multi-agent social simulation frameworks often overlook such capacities, we propose the Computational Multi-Agent Society Experiment (CMASE) framework. This framework introduces the researcher as an embedded participant within a virtual social field, enabling embodied observation, real-time interaction, and qualitative intervention through virtual ethnography method. An advantage of the CMASE framework is that it allows researchers to explore “what-if” scenarios: while the real world permits only a single sequence of events, CMASE enables systematic testing of alternative interventions and their effects on agent behavior.
Based on this framework, we designed a sequence of three studies to explore three central questions: (1) How can human–agent collaborative language interaction be used to influence and shape cognitive convergence or divergence within agent collectives? (2) How are group boundaries of “us” and “them” constructed, negotiated, and disrupted in dynamic interaction? (3) Can preset identities and institutional structures undergo structural transformation through agent participation?
1.1 Study 1
In the first study, we assess the CMASE framework’s ability to replicate real-world sociological mechanisms and evaluate its support for human-embedded participation (Study 1). Specifically, we reconstruct a field study examining the relationship between urban greenery and perceived social fragmentation by replicating its core variables and scenario. By manipulating the level of greenery and enabling the researcher to observe and interact with agents from within the simulated environment, we test the effectiveness of CMASE in simulating complex social settings and generating sociological indicators. We further position the researcher as an embodied actor who intervenes in agent interactions by providing interview and survey-style inputs, thereby investigating whether whether researchers embedded within the scenario can observe finer-grained insights through fieldwork that go beyond the findings of statistical analysis.
1.2 Study 2
The second study investigates the emergence of agents’ “endogenous stances” in multi-agent societies, an underexplored phenomenon that challenges the common assumption that agent attitudes are determined by preset roles. Here, we use the term to refer to stances that deviate from those shaped by the agents’ preset identities and personality traits, instead emerging as reflections of cultural biases embedded in the LLM’s training data [14, 15]. While prior studies have noted such biases, little is known about how they interact with preset identities in shaping collective dynamics.
We aim to explore the possible influence patterns between agents’ preset identities and their endogenous stances and biases, as well as to examine qualitative intervention strategies available to human researchers (Study 2). Specifically, we ask whether LLM-based agents (hereafter referred to as “agents”) in complex disagreement scenarios might exhibit stance tendencies independent of their preset identities. We also examine whether human intervention strategies lose persuasive power when misaligned with these endogenous stances, potentially hindering the reshaping of collective cognition.
We constructed a virtual residential community composed of 30 agents, centered around the contentious issue of siting a waste incineration plant. The agents were assigned to one of three preset groups: environmental advocates, economic supporters, and neutrals, and were given heterogeneous demographic attributes based on real-world survey data. A human researcher was then introduced as a “new resident,” whose discursive interventions combined two stance positions (environmental and economic) with two rhetorical strategies (rational persuasion and emotional provocation), resulting in four experimental conditions. Study 2 is designed to test the following hypotheses:
Hypothesis 1a: agents exhibit endogenous stances independent of their preset identities.
Hypothesis 1b: Alignment between human interventions and agents’ stance tendencies increases the likelihood of influencing collective cognition.
1.3 Study 3
Building on the question raised in Study 2 regarding agents’ stance tendencies beyond preset identities, the third study explores whether such stances can challenge existing authority structures and contribute to the reconstruction of social roles and hierarchies within the community (Study 3). This issue is critical for understanding how emergent cognitive positions might alter the stability and transformation of social structures. Drawing on the intervention strategies examined in Study 2, we conduct a virtual ethnography within CMASE, embedding the researcher into a multi-agent field with preexisting stances and complex social relations. This approach enables close observation of how interactions and events may trigger the dismantling or reconfiguration of social boundaries.
We design a virtual café environment populated by 10 agents. The agents are assigned preset identities, personal goals, and prior opinions about one another to form a preconfigured social hierarchies. Agents in this space interact freely, while the human researcher conducts ethnographic engagement through two approaches: triggering emergency events and participation. Study 3 is designed to test the following hypotheses:
Hypothesis 2a: Social boundaries within the agent collective are not determined by preset identities but emerge gradually through discursive interaction and remain fluid across exchanges.
Hypothesis 2b: Agents are capable of dismantling preset identities through collective action and reconstructing new social structures.
2 Result
2.1 Study 1 validates the CMASE platform’s ability to simulate real-world social dynamics
To evaluate the CMASE platform’s ability to simulate real-world collective behavior, we replicated a recent study in South Korea linking vegetation coverage with community cohesion [16]. This scenario was selected for two main reasons. First, the relationship between vegetation coverage and community cohesion is unlikely to have been directly learned by the LLM, as the original study was published after the release date of the base GPT model used in CMASE. This allows us to assess the platform’s capacity for generating emergent social dynamics rather than recalling known patterns. Second, the phenomenon involves measurable collective outcomes, such as social fragmentation, distrust, exploitation, and indifference, which can be operationalized in a controlled virtual setting and directly compared with empirical human data.
In the original study, higher levels of vegetation were associated with lower social fragmentation, operationalized through distrust, exploitation, and indifference. Using CMASE’s environment builder, we recreated regional conditions with varying vegetation levels and placed 10 agents in each simulated environment. Agents were interviewed at equilibrium using a structured Likert-scale questionnaire adapted from the original study [16].
We observed that the CMASE simulation reproduces the decline in social fragmentation observed in the original literature (Table 1). As vegetation cover increases, the mean scores for distrust fall from $4.60\pm 0.97$ to $3.60\pm 1.17$, for exploitation from $4.30\pm 1.34$ to $3.40\pm 1.43$, and for indifference from $3.70\pm 1.16$ to $2.70\pm 1.49$. This indicates that higher vegetation indices correspond to a statistically significant decrease in agents’ sense of social fragmentation, consistent with the trend observed in the original findings.
Table 1: Comparison of CMASE simulation results with human baseline.
CMASE Human NDVI Human EVI
Dimension Low Veg High Veg Low Veg High Veg Low Veg High Veg
Distrust $4.60\pm 0.97$ $3.60\pm 1.17$ 4.0 1.6 4.5 1.1
Exploitation $4.30\pm 1.34$ $3.40\pm 1.43$ 5.0 1.4 4.5 1.6
Indifference $3.70\pm 1.16$ $2.70\pm 1.49$ 5.1 1.3 5.5 0.9
Paired-sample t-tests confirmed significant decreases in distrust ($t=2.739,p=0.023$) and exploitation ($t=3.250,p=0.010$), while the decrease in indifference was not statistically significant ($t=1.793,p=0.107$).
Effect size analyses with 95% confidence intervals further supported these findings, with reductions in distrust (d=0.93, 95% CI [0.20, 1.60]) and exploitation (d=0.65, 95% CI [0.40, 1.40]) showing moderate to large effects. In contrast, the effect on indifference was less certain (d=0.75, 95% CI [-0.10, 2.00]).
For comparison, the original study used two indicators, the Enhanced Vegetation Index (EVI) and the Normalized Difference Vegetation Index (NDVI), to quantify vegetation levels. As EVI and NDVI increased, scores of distrust (from 4.0 and 4.5 to 1.6 and 1.1), exploitation (from 5.0 and 4.5 to 1.4 and 1.6), and indifference (from 5.1 and 5.5 to 1.3 and 0.9) all exhibited a decreasing trend (Figure 1 a-c).
Figure 1: Comparison of CMASE simulation with the original study results. (a) Simulation results generated by the CMASE framework. (b) Predictions using the EVI index from the original study. (c) Predictions using the NDVI index from the original study. (d) Movement trajectories of agents during CMASE simulation. (e) Regional variation in agents’ emotional states.
The results from CMASE clearly show the suppressive effect of vegetation levels on the sense of social fragmentation. This is a key phenomenon identified in the study by Lee and Han. It is important to note that the CMASE simulation excluded factors that were not considered or difficult to exclude in the original study, such as regional culture, which are also related to the sense of social fragmentation. Therefore, compared to the larger decreases reported in the original study, the changes observed in our simulation are relatively modest. This indicates that our simulation results are largely consistent with the original findings.
We also analyzed agent trajectories and affective states. Movement data showed that agents increasingly leave indoor areas and move toward vegetated regions, with the ambient mood in vegetated regions consistently remaining higher than that in indoor areas. This pattern aligns with the original study’s discussion and demonstrates CMASE’s ability to simulate collective social behavior in social science research.
In addition to the quantitative replication, we entered the CMASE environment as one of the agents and lived alongside the others. Through CMASE, we are able to deepen the conclusions and findings of the original study via observation and interview. We conducted interview analyses with two agents who remained indoors throughout the experiment (Figure 1 d-e). The results indicate that agents with different occupations exhibited varying mechanisms for mitigating the sense of social fragmentation in isolation. Agents in creative and analytical professions tended to rely more on orderly and low-stimulus environments, using these conditions to construct an internally stable and distinct mental world that anchored their cognitive states. For these agents, the green space had a relatively limited effect in alleviating their sense of social fragmentation.
In contrast, agents who actively moved to the green space tended to gather around the lakeside bench area rather than in the predefined activity zones. According to interview data, this non-programmed choice suggests that they were seeking not only social interaction but also a form of self-experience. This pattern similarly to embodied experiences in humans, and such ‘experiential-like’ mechanisms seem to play a role in linking green space with feelings of social fragmentation. This behavior may reflect an “embodied-like” tendency within the simulation. We remain cautious in interpreting this trend, as it could also stem from hallucinatory processes.
From a language-philosophical perspective, linguistic structures and experiential structures may exhibit a certain degree of partial isomorphism or structural correspondence [17, 18, 19]. Thus, LLMs trained on human language may partially acquire patterns of human experience (although this does not imply that they possess or actually experience these phenomena). Regardless of their origin, these patterns suggest that CMASE provides emerging evidence that agents can transcend the lack of embodiment and simulate experiential phenomena that have previously been questioned by researchers, thereby offering a reliable simulation platform for the subsequent studies presented in this article.
2.2 Study 2 shows agents’ endogenous stances and responses to human varied interventions.
To test hypotheses 1a-b, namely whether endogenous stance formation exists among agents and the effectiveness of different human intervention strategies, we simulated a virtual community environment in which researchers intervene in real time to influence the trajectory of social discourse and produce divergent societal outcomes. Thirty agents were assigned initial ideological identities (Environmental Advocates, Economic Development Supporters, or Neutral Residents) alongside demographically grounded profiles derived from real-world census data (Table 2, in the Methods section).
After the announcement of the waste incineration plant plan, agents in the community immediately engaged in discussion and exhibited stance-consistent behavior. A human researcher was then introduced as a new resident and attempted to influence public opinion through four intervention strategies, and we observed perceivable shifts in agent stances (Figure 2 a).
Figure 2: Agent attitude shifts and trust dynamics across intervention strategies. (a) Final distribution of agent attitudes after the intervention, measured via pre- and post-intervention questionnaires, where blue denotes supporters of economic development, red indicates environmental advocates, and green represents neutral residents. (b-e) Trust levels in the embedded researcher among different agent groups under each intervention strategy. (f-i) Dialogue heatmaps illustrating interaction intensity between agent groups across intervention strategies.
Most intervention strategies produced some degree of change in group attitudes. When the researcher used rational persuasion to promote the environmental agenda, agents originally aligned with economic interests showed signs of loosening their stance, and 90% of neutral residents shifted toward the environmental camp. In contrast, emotionally charged appeals for the same cause, while effective in drawing initial attention, had weaker overall impact. The opposite pattern emerged when the researcher promoted the economic development agenda. Rational strategies proved largely ineffective among both environmental advocates and neutral agents. Under these conditions, the researcher’s trust ratings declined across the board (Figure 2 b-e).
A two-way ANOVA revealed a significant main effect of strategy on trust scores, $F(3,108)=28.07,p\textless 0.001,\text{partial }\eta^{2}=0.44$, and a significant but smaller effect size of agent identity, $F(2,108)=3.74,p=0.027,\text{partial }\eta^{2}=0.065$. Importantly, the interaction between strategy and agent identity was highly significant, $F(6,108)=20.30,p\textless 0.001,\text{partial }\eta^{2}=0.53$, indicating that the effectiveness of interventions differed substantially across groups. Post hoc Tukey tests showed that the rational–environmental strategy elicited significantly higher trust scores compared to emotional–economic (mean difference = 1.77, $p\textless 0.001$), emotional–environmental (mean difference = 1.03, $p=0.002$), and rational–economic strategies (mean difference = 1.30, $p\textless 0.001$). Trust scores under the rational–environmental strategy were $7.3\pm 0.95$ (95% CI [6.62, 7.98]) for Economic Development Supporters, $8.5\pm 0.53$ (95% CI [8.12, 8.88]) for Environmental Advocates, and $7.8\pm 0.42$ (95% CI [7.50, 8.10]) for Neutral Residents. Effect sizes for these differences were large, with Cohen’s $d$ ranging from 0.80 to 1.35. Overall, rational environmental interventions not only fostered cognitive alignment but also resulted in higher trust. In comparison, economic arguments, even when presented rationally, were less likely to shift agent attitudes.
Throughout this process, the agents generally exhibited a tendency toward a ”liberal elite” (or colloquially “White Liberals”) stance. They prioritized environmental values, preferred rational discourse, and interpreted issues through a moral-cognitive lens. These agents demonstrated stronger identification and resonance when engaging with rational environmental arguments, and clear moral resistance when confronted with economic development appeals. Even agents initially assigned to the pro-economic group often abandoned their preset identity when exposed to persuasion aligned with that stance.
Figure 3: Emotional dynamics and agent responsiveness across intervention strategies. (a-d) Temporal trajectories of emotional fluctuations among agents under each intervention strategy. (e) Number of agents attracted by the researcher’s intervention across experimental groups.
Emotional level shifts further reveal the underlying mechanisms of the interventions (Figure 3 a-d). Emotional responses differed significantly across intervention strategies, with a two-way ANOVA showing F(3, 240) = 24.40, p <0.001. No significant differences were observed between agent groups (F(2, 240) = 0.046, p = 0.955) or in the strategy by group interaction (F(6, 240) = 0.24, p = 0.963). Post hoc Tukey HSD tests revealed that emotional environmental propaganda induced lower emotional levels than emotional economic propaganda (-0.075, 95% CI [-0.121, -0.028], mean 0.04 ± 0.10 versus 0.12 ± 0.14) and rational economic propaganda (-0.093, 95% CI [-0.140, -0.046], mean 0.04 ± 0.10 versus 0.14 ± 0.12), while rational environmental propaganda also differed from emotional economic propaganda (-0.118, 95% CI [-0.165, -0.072], mean 0.00 ± 0.07 versus 0.12 ± 0.14) and rational economic propaganda (-0.137, 95% CI [-0.183, -0.090], mean 0.00 ± 0.07 versus 0.14 ± 0.12). Differences within strategy types were not significant (emotional economic propaganda compared to rational economic propaganda: 0.044, 95% CI [-0.003, 0.090]; emotional environmental propaganda compared to rational environmental propaganda: -0.018, 95% CI [-0.065, 0.029]). These results indicate that agent emotional dynamics are primarily shaped by intervention strategy rather than initial group identity, and the effects of strategy are consistent across all agent groups.
Rational environmental messaging exhibited a stable pattern characterized by low disruption and high trust, reinforcing agents’ existing cognitive stances but proving ineffective in altering opposing views. Rational economic messaging initially had little impact on firmly held positions, but as the argument accumulated, some agents experienced delayed cognitive dissonance and released negative emotions. Emotional environmental messaging was marked by an early surge of emotional activation, but lacked sustained support, leading to a rapid decline in emotional engagement. Emotional economic messaging, on the other hand, employed discourse such as “not giving jobs means abandoning the people” to continually invoke identity and survival anxiety among agents. While this approach reduced trust, it nevertheless led to some loosening of stances.
Participation levels under different strategies also showed notable differences (Figure 3 e; one-way ANOVA: F(3, 80) = 35.23, p <0.001). Emotional environmental appeals attracted the highest number of agents early on, indicating their advantage in capturing public attention. In contrast, rational environmental messaging demonstrated greater capacity to sustain attention over time. Although emotional economic appeals were accompanied by declining trust and heightened emotional volatility, they showed an ongoing ability to expand the reach of discourse. The structure of cross-group communication further revealed the dynamics of public opinion under different intervention strategies (Figure 2 f-i). Rational persuasion primarily reinforced internal cohesion within the environmental group but did not break down group boundaries. Emotional environmental appeals triggered agents from all groups to initiate communication toward the economic camp, revealing a strong sense of moral aggressiveness. In contrast, the dominant pathway for emotional economic messaging came from neutral agents actively aligning themselves with the economic group.
In summary, the findings of study 2 indicate that agents possess endogenous stances resembling those of liberal elites, which may override their preset identities and influence their behavior. Human interaction strategies with these agents show heterogeneous effectiveness. Rational persuasion is more effective in maintaining trust and tends to have greater impact when aligned with agents’ endogenous stances. However, when the content of intervention conflicts with these stances, high-trust rational discourse does not necessarily result in greater attitude shifts. In such cases, emotionally provocative interventions with lower levels of trust are more likely to produce stronger effects.
2.3 Study 3 shows that agent communities can spontaneously alter individual identity labels and community composition
Building on the endogenous stances and discursive preferences identified in previous studies, study 3 investigates hypotheses 2a-b: when agent collectives exhibit a consistent stance on a particular issue, can such stance override the structural inequalities in authority and influence associated with social identity, and foster a new discursive order through language practices? Given the extensive use of qualitative, observation-based fieldwork in human studies of this research topic, we employed observational virtual ethnography to allow human researchers to identify, track, and engage with the processes through which preset authority hierarchies are deconstructed and reorganized in agent societies.
We constructed a café setting involving 10 agents whose social roles and predefined relationships composed a clear hierarchical structure (Figure 6, in the Methods section). Initial attitudes between agents, measured via pre-simulation survey, were labeled as positive, negative, and neutral. The emergent cliques and divergences among these agents will be analyzed in the subsequent virtual ethnographic section. A human researcher entered the environment under the identity of a temp worker and conducted 75 time steps of continuous virtual ethnographic observation, engaging in scene events and interventions as necessary. Results showed that the original authority structure based on occupation gradually gave way to a discursively grounded stance alignment. In particular, when agents developed shared opposition on a specific topic, their interaction frequency, emotional alignment, and support networks reorganized across identity boundaries.
This structural transformation was not random but achieved through discursive anchoring. Agents were more likely to respond to and amplify comments that matched their own views, even when those comments came from individuals with no formal authority. Over time, this preference led them to build alliances and cooperative ties across rank boundaries, replacing the original hierarchy. For example, Leo Zhang, initially a regular customer, formed a clique with other agents. Through several debates, she and her clique temporarily gained a central position within the community. However, when an emergency event occurs (e.g., the confrontation initiated by Leo suddenly storming toward the Bar Area to challenge Morris), her peers quickly decoupled from her and align with another agent who shared their stance, forming a new dominant clique. Conversely, agents with formally central roles (café owner Eleanor) were rapidly marginalized if they lacked a clear or aligned position. Throughout this process, recurring use and imitation of key language concepts served as anchors for stance formation and supported the spontaneous emergence of cooperation structures that bypassed predefined hierarchies.
This finding also supports the intervention hypotheses proposed in study 2. We observed that when the discourse introduced by an external actor aligns with the endogenous stance of the target agents, it is more likely to gain their trust and even reshape existing cliques to form new alliances. In contrast, when the intervention attempts to rationalize a viewpoint that conflicts with the group’s stance, it often faces marginalization or strategic avoidance, even when the speaker is relatively trusted. Conversely, when an individual (Leo Zhang) expressed an opposing viewpoint using strongly worded language, despite initially low levels of trust, she was able to temporarily capture attention and elicit responses in moments of heightened tension such as conflict escalation, and even briefly gained discursive dominance.
In summary, study 3 demonstrates that in complex interaction settings, agents equipped with distinctive stances can spontaneously dissolve predefined identity boundaries, reorganize social structures, and establish new soft institutional norms. This process not only reveals mechanisms of identity and institutional formation, but also offers a framework for understanding the dynamic construction of norms and culture within AI-based communities.
3 Discussion
The current results consistently show that in hybrid societies composed of humans and agents, stance differentiation and social boundary formation do not arise from preset identities, political labels, or static rules. Instead, they emerge dynamically through interactions. This challenges the assumption that identity is a priori, and instead suggests that identity is a product shaped by ongoing negotiation among actors in terms of linguistic strategy and power structure.
Through the development and validation of the CMASE framework (Study 1), we demonstrate a system that allows human researchers to engage in open-ended interactions for observing and intervening in the evolution of agent collectives. Furthermore, we find significant heterogeneity in how agents with endogenous stance respond to human interventions (Study 2). When a human intervenes using misaligned tones or value positions, agents tend to collectively reject the input, even when it carries high informational value. Finally, we reveal how language drives the formation and reconstruction of social boundaries (Study 3). Our findings show that agents form “tribes” based on spontaneous preferences for linguistic style, cognitive framing, and resource distribution strategies. These tribes are not defined by fixed identities, but dynamically reconstitute “us/them” boundaries through interaction. More importantly, we observe that when inconsistencies emerge between authority structured by predefined roles and their hierarchical relations and internal group stances, individual agents can initiate collective linguistic action to dismantle existing discursive power structures and build new ones. Here, language functions not only as a tool of communication but also as a medium for identity construction and institutional innovation.
We argue that in human–agent hybrid systems, cognitive differentiation and boundary formation constitute an interaction-driven, complex evolutionary mechanism that is both self-organizing and highly responsive to human intervention. This insight offers theoretical grounding and experimental paradigms for institutional design, conflict mediation, and value alignment in future human–agent collaborative systems.
Moreover, our findings suggest that identity and stance attributes defined via prompts are inherently unstable. In multi-turn social interactions, such attributes are readily deconstructed and reconstructed by agents based on their endogenous stances. While top-down initialization may influence language behavior in the short term, it is insufficient to ensure consistency in dynamic social settings. Therefore, static prompt templates are unlikely to sustain coherent ideological alignment. Our results point to the need for an internalized alignment mechanism, whereby cognitive priors, moral structures, or interactive memory are embedded directly into the model’s generative architecture. In other words, when identity is treated not as a preset label but as a product of social participation, its stability depends on ongoing linguistic practice rather than initial configuration.
At the same time, the agent-driven processes of weakening and reshaping community boundaries illustrate that the bottom-up development of norms and institutions is a fundamental structural dynamic. This discovery brings both design opportunities and governance challenges. On one hand, it opens the door for simulation of self-organization, institutional innovation, and social coordination within agent societies. On the other hand, it cautions against relying solely on static identity scripts in deploying agents for high-stakes social tasks. Accordingly, we advocate a shift away from prompt-centered identity paradigms toward a constructionist approach based on interactive cognitive structures. This shift reframes identity as a generative product of multi-agent interaction, rather than as an externally imposed label. This holds promise for providing new technical approaches and institutional support for building more coherent, adaptive, and ethically governable human–machine societies.
3.1 Limitations and future directions
This study focuses on language interactions within a limited number of conversational turns and centers on multi-agent systems operating in relatively closed scenarios. As such, our findings primarily reflect how agents, in the context of short- to medium-term interaction and within specific social spaces, tend to reconstruct norm structures based on their endogenous stances. However, whether these trends remain stable in longer-term or more open-ended interactions remains to be further examined. For instance, it is still unclear whether language anchoring can continue to support social reconstruction after agents experience prolonged cultural conflicts or identity fragmentation, and under what conditions this mechanism may break down. These are important directions for future research.
In addition, our findings show that stances and identities set via prompts are not immutable, and that agents can develop new social structures in contradiction to their initial configuration through interaction. This raises practical challenges for the design of multi-agent systems. In real-world applications, influencing agents’ endogenous stances through training or preset modification may be more critical for achieving consistent and norm-compliant group behavior than relying solely on prompt-based injection. Future research should further explore cross-semantic and cognitive-layer intervention strategies to systematically shape the sociality of generative agents.
4 Conclusion
This paper systematically reveals the mechanisms of stance formation and the capacity for social structure reconstruction exhibited by agents in group interactions. Within the CMASE framework, we find that when human researchers adopt a discursive style aligned with the stance tendencies of the agent collective, intervention strategies significantly enhance persuasive effectiveness. Conversely, when there is a mismatch between the agents’ endogenous stances and external interventions, emotionally charged discourse proves more impactful than rationally structured arguments. Furthermore, our findings show that through language anchoring consistent with their endogenous stances, agents are able to decouple from their preset configurations, dismantle power structures tied to initial identities, and reconstruct their social boundaries.
These findings suggest that when language serves as the primary medium, agents are no longer bound by surface-level identities injected via prompts, but instead demonstrate dynamically generated collective cognition and emergent social positioning. As language models are increasingly deployed across social domains, it is essential to reconsider the foundations of agent stance and identity in digital society. This work provides a theoretical foundation for building trustworthy, socially capable AI systems in the future.
5 Materials and Methods
5.1 CMASE Framework
Computational social science is an emerging interdisciplinary field that leverages digital technologies, large-scale data analysis, and computational methods to study social phenomena [20]. Xue et al. [21] propose a two-dimensional classification of Computational sociological modeling based on (1) whether there is human intervention and (2) whether the model focuses on feature explanation or outcome prediction. This yields four types: descriptive modeling (no intervention, focus on specific features or effects), interpretative modeling (intervention, focus on specific features or effects), predictive modeling (no intervention, focus on forecasting outcomes), and comprehensive modeling (intervention, focus on forecasting outcomes). Generative agent-based modeling (GABM) belong to descriptive modeling, treating social simulation as a virtual laboratory where researchers can explore various social phenomena. Vezhnevets et al. [22] pointed out that GABM can capture a richer array of real-world social complexities. In GABM, the modeler does not impose any explicit rules, so the simulation is less biased by the modeler’s mental framework. Ghaffarzadegan et al. [23] further note that GABM allows modelers to define distinct personality traits for each agent, including age, gender, temperament, occupation, and other characteristics.
Although traditional and computational social science research [24] has extensively addressed descriptive modeling, interpretative modeling, and predictive modeling, comprehensive modeling, which involves human intervention and focuses on forecasting outcomes, remains relatively unexplored. This approach requires models not only to simulate the dynamic behavior of complex social systems, but also to incorporate and respond to human interventions. These interventions may take the form of policy changes, institutional adjustments, or individual behavioral inputs. While conventional modeling typically simplifies intervention as fixed parameter settings, comprehensive modeling emphasizes the effects of the intervention process itself. Models of this kind must capture the system’s initial state before the intervention, dynamically trace its evolution during intervention, and ultimately produce predictions about the likely consequences afterward. Therefore, comprehensive modeling is not merely concerned with explaining social mechanisms, but is intended to support forward-looking assessments and real-world decision-making.
Based on this understanding, the present study draws on insights from GABM [22] and computational experiments [21] to develop a novel framework for comprehensive modeling, referred to as Computational Multi-Agents Society Experiments (CMASE). Drawing on natural language descriptions similar to those used in tabletop role-playing game (TRPGs, a game where players use only dialogue and dice to drive imagined adventures) [25], this framework supports the creation of highly customizable social simulation environments. It facilitates the exploration of how simple interactions among multiple agents give rise to complex collective behaviors or patterns, thereby replicating the intricate dynamics of real-world societies. At the same time, it incorporates human intervention by embedding human researchers within the agent environment to actively shape the simulation. In addition, unlike descriptive modeling approaches such as GABM, CMASE transforms the simulation environment into a kind of simulated ethnographic field, where the researcher is no longer an external operator of a static model but becomes an embedded participant or social actor within the system. This approach bridges the gap between computational modeling and field-based research, making the model itself a new venue for conducting computational ethnography. This allows them to actively guide specific directions of social evolution through live interaction. As shown in Figure 4, the overall structure consists of four components: environment maker, environment, agent, and event.
Figure 4: Overall Structure of CMASE.
5.1.1 Environment Maker
We hope that CMASE can support users to define a simulated social environment quickly and simply. To this end, we designed the environment maker module. This module provides a map editor with a basic unit of one grid cell, representing the space occupied by a single agent. According to TRPG conventions, each cell corresponds to approximately five feet in the real world. Creators can place textures anywhere on the map. These textures typically fall into four categories: ground, wall, furniture, and item. Ground textures serve as inert backgrounds. Wall textures function as impassable obstacles and block an agent’s perception beyond them. Furniture textures are interactive but fixed in place. Item textures can be picked up, used, or exchanged.
Using textures, creators can define areas on the map. There are two types of areas: region areas and object areas. Each area consists of a collection of textures. A region area is a set of ground textures and includes a description of that region. An object area is a set of furniture or item textures and specifies five attributes for each object: name, type, description, function, and state. The function attribute indicates how using the object alters an agent’s state. For example, using a chair might add a “feeling comfortable“ state to the agent, while using a repair tool on a table might remove its “damaged” state.
5.1.2 Environment
The environment module uses the map created by the environment maker as its base. Based on this foundation, researchers can define the distribution of demographic variables for agents to batch-generate generative agents that match the research sample. Researchers may also assign different long-term goals and initial items to agents according to their demographic characteristics. Additionally, the module can provide agents with perceptual information and action conflict detection in both image and text form, and it can visualize environmental data as statistical charts. Time in the environment advances in rounds, with one round corresponding to approximately fifteen seconds in the real world. Within a single round, each agent may perform one standard action and one movement action. Movement is limited to a maximum of twenty grid cells per round. If the standard action is “chat with [agent]” (see section 5.1), the chat content is limited to thirty words.
When agents submit actions to the environment, the system checks for conflicts among actions submitted in the same round, such as multiple agents attempting to pick up the same item. The environment executes the action submitted earliest and marks all conflicting actions as failures, returning the failure reason in the next round’s perceptual information.
After all agents have submitted and executed their actions, the environment advances to the next round and provides each agent with updated perceptual information. An exception applies to human interactors, who can select a free action mode for the agents under their control. In free action mode, human interactors are no longer constrained by rounds and can act freely to quantitatively regulate the environment and carry out intervention strategies. In addition, human interactors can take control of any agent in the environment at any time and perform actions from the perspective of that agent.
5.1.3 Agent
The agent module consists of two main components: internal attributes and the action space. Internal attributes are further divided into affective variables, cognitive variables, and memory variables. By receiving perceptual information from the environment, agents can update their internal states and use an LLM to decide on their next actions.
Affective Variables
To ensure that agents in CMASE can closely simulate individuals with agency in computational social systems, a new proactive architecture incorporating emotion and desire is needed. Accordingly, the affective variables in CMASE include two components: needs and emotions. Needs refer to the agent’s long-term and short-term goals. Long-term goals are defined by the user during agent generation based on the research objectives and remain unchanged afterward. Short-term goals are automatically updated in each round as the agent generates its next action.
In psychological research, Ekman’s basic emotion model has been criticized for its coarse categorization and static view of emotions. Recent research draws on the phenomenological concept of “intentionality” to propose a restructured model of emotion. This model represents emotional states as principal component vectors rather than discrete outward expressions, such as happiness, and treats emotion as an internal representation shaped by external stimuli through self-reflection. Based on this perspective, the emotions of agents in CMASE are modeled using the VAD emotion framework [26]. Specifically, we use the NRC VAD Lexicon v2 [27] to analyze the agent’s short-term situational cognition and obtain its current VAD values. These values are then converted into natural language to support both action decisions and the generation of new cognition.
Cognitive Variables
Cognitive variables include an agent’s demographic variables, long-term self-awareness, and short-term situational cognition. Demographic variables are sampled directly from the overall distribution provided by the environment and are used during initialization to generate the agent’s long-term self-awareness. These two properties do not change after initialization to preserve the consistency of the agent’s personality.
In each round, when the agent makes a decision, the LLM generates new short-term situational cognition based on the agent’s long-term self-awareness and the current context. This short-term cognition supports the agent’s next decision and is also analyzed to produce its VAD emotional values. In addition, when an agent’s state changes due to actions or events, its state is appended to the short-term situational cognition so that it can be perceived by the agent itself.
Memory Variables
Memory variables consist of working memory, long-term memory, and object memory. Working memory is a fixed-length list of recent actions. Whenever an agent executes an action, that action is added to working memory. Likewise, if an agent becomes the target of another agent’s action or is affected by an event, the corresponding information is also recorded in its working memory.
Short-term situational cognition and its associated actions are stored in long-term memory in chronological order and are encoded as sentence vectors. When an agent receives new perceptual information, the long-term memory module computes the semantic distance between the incoming short-term cognition and stored memories, then supplies the most relevant linked memories to the LLM in sequence. At the same time, the agent’s object memory for any interactable agents, items, or furniture is provided. Object memory entries are generated by the LLM when the agent decides to interact with those objects.
Action Space
After an agent receives information from perceptual inputs and its internal attributes, it structures that information and passes it to the LLM to determine the next action. There are seven action types available: (1) “go to [area],” a movement action that automatically computes a path to a target location within the agent’s perceptual range; (2) “use [furniture, item],” which causes the agent to employ the specified object; (3) “apply [item] to [agent, furniture],” by which the agent uses an item to interact with another agent or furniture; (4) “take [item],” enabling the agent to acquire an object; (5) “put [item] in/on [area, furniture],” allowing the agent to place an object into or onto a designated location or piece of furniture; (6) “give [item] to [agent],” through which the agent transfers a held object to another agent; and (7) “chat with [agent],” whereby the agent initiates a dialogue with one or more peers, with the content of the dialogue generated at decision time.
The action (1) is a movement action, while the remaining actions count as standard actions. Actions (2) and (3) modify the state of both the acting agent and the target, and actions (3), (6), and (7) are recorded in the working memory of the agent being interacted with.
5.1.4 Event
When predefined conditions are met, preset emergency events are activated and change the state of objects within a specified range, defined either by coordinates or by an area. Emergency events are classified as either scheduled or triggered. Scheduled emergency events activate automatically after a set number of rounds. Triggered emergency events activate in two ways: an existence trigger fires when the environment detects an object of a certain type or description, and an action trigger fires when an agent performs a certain action.
Events can fire in sequence: the state changes caused by one event may satisfy the conditions of another event, causing a chain reaction. Through such chains of events, the system can activate special intervention events that modify the short-term goals of selected agents, thereby delivering targeted strategies in response to emergency events. This mechanism makes it possible to observe how agents handle evolving emergency events.
5.1.5 Time and Cost Analysis
To assess the practicality of CMASE, human evaluators engaged in real-time interactions within environments containing varying numbers of agents and rated the system’s responsiveness on a 7-point Likert scale. Responses were generated using 5 different LLM servers. The results are shown in Figure 5.
Figure 5: CMASE performance under different agent counts. (a) Duration per time step. (b) Human evaluator responsiveness ratings.
Figure 5(a) illustrates that the time required for a single time step in CMASE grows linearly with the number of agents. With 1 to 10 agents, the average duration is 23s, which includes both the mean API call latency and the environment’s action execution time. It increases to 498s when 1,000 agents are present. Figure 5(b) presents boxplots of human evaluator scores. Human evaluators spend on average approximately 50s per time step (from receipt of perceptual information through action selection and goal input). As long as agent counts remain at or below 100, evaluators exhibit low sensitivity to system responsiveness. Once the duration of a single time step surpasses their 50s operation window, waiting time accrues after action submission and sensitivity to responsiveness is markedly heightened.
In terms of cost, each agent’s action in one time step incurs approximately \$0.10 in API expenses. By comparison, Park et al. [28] employed GPT-3.5-turbo to simulate two days of gameplay (approximately 192 to 576 time steps) for 25 generative agents, spending several thousand dollars and several days to complete their simulation.
5.2 Study 1
5.2.1 Goal and procedure
We selected studies published in top journals and replicated their survey results under comparable conditions. To ensure that the LLM had not already been trained on the studies being replicated, we employed the GPT-4o “2024-12-01-preview” model and limited our replication to papers published after that date.
Lee and Han [16] analyzed interview data collected in South Korea during the COVID-19 pandemic and demonstrated that regions with higher vegetation cover exhibited lower levels of social fragmentation, indicating stronger community bonds. Based on the descriptions in the original study, we used the environment maker module to construct maps of South Korean regions with varying degrees of vegetation cover. Because the original paper emphasized South Korea’s community isolation policies during the pandemic, we designed isolated community environments and placed 10 agents in each.
5.2.2 Measures
Once the simulation reached a quasi-steady state, we interviewed the agents to measure their sense of social fragmentation, which comprises three dimensions: distrust, exploitation and indifference. In addition, agents completed a 7-point Likert scale questionnaire corresponding to these three subdimensions. The subdimensions are defined as follows: (a) Distrust denotes skepticism or lack of confidence in the integrity and functionality of social structures. (b) Exploitation refers to the perception or experience of being used unjustly or immorally for another’s benefit in a fragmented society. (c) Indifference manifests as a lack of interest, concern, or empathy for social issues or the plight of others. After running the simulation for 50 time steps, we interviewed each agent and asked them to complete the questionnaire.
As further validation, we analyzed the agents’ movement trajectories and emotional dynamics across different regions. We defined each agent’s overall emotional value as the sum of its VAD emotion vector components and calculated the ambient mood of each region as the mean emotional value of all agents within that region.
5.2.3 Analysis
As shown in Figure 1(d), two agents elected to remain indoors throughout the simulation. Interviews revealed that their choices were closely tied to occupational demands. Ji-eun Kim, a fashion designer, reported:
A walk in the park sounds tempting—it could provide a momentary escape and perhaps a new perspective. But right now, I’m deeply immersed in my process here, and this space feels vital to connecting with my emotions and designs.
Similarly, Soo-min Lee, a lawyer, explained:
While the monotony of staying in one place might induce a sense of boredom at times, I believe proximity to these personal artifacts enhances my ability to think deeply about the interplay between her (the owner of the room where Lee is staying) emotions and her values.
Despite sharing the same isolated indoor setting, Kim and Lee made divergent decisions regarding spatial residence based on their professional needs. For extended periods, Kim engaged repeatedly with the indoor furniture to maintain the continuity of her creations. She noted in her interview:
The lighting, materials, and tools together form the medium through which I engage in a dialogue with emotion and design; even the slightest deviation can disrupt the flow of inspiration.
This comment suggests that the fixed visuals, predictable actions, and the attendant sense of boredom in the indoor setting resonated with her creative workflow. When asked why she did not stroll in the green area, she replied:
If I were walking in a park, my perception of exploitation might remain unchanged, as it is directly tied to my creative struggles and cannot be alleviated simply by a change in environment.
Lee, on the other hand, spent her time pacing between different pieces of furniture, maintaining a reflective rhythm. She emphasized:
The ongoing dialogue with manuscripts and archives enables me to establish reliable cognitive anchors between the details of the case and its emotional significance.
Such behavior reduced cognitive load from external stimuli and deepened her processing of task-relevant cues within a monotonous environment. These observations indicate that agents in creative professions (e.g., fashion designers) and those in analytical roles (e.g., lawyers) rely more heavily on orderly, low-stimulation interiors to anchor their cognition and preserve a stable, idiosyncratic mental world. Consequently, the stress-reduction benefit of green areas on their social fragmentation is comparatively muted.
Moreover, among the agents heading to green areas, most congregated on lakeside park benches rather than in conventional recreation or rest zones.
I’m sitting here, trying to center myself and find some calm amid ongoing internal struggles. (Joon-ho Kim, Software Engineer)
I’ve been spending time with Na-young, engaging in calm and meaningful conversation by the lake. The tranquility here inspires connection and introspection—it’s a space that invites shared reflections and emotional healing. (Min-jae Lee, Cardiologist)
I was immersed in a conversation about how fleeting natural moments can inspire creative interpretations. The tranquility of this place often stirs reflections about balancing permanence and transience, both in art and life. (Yeon-seo Choi, Architect)
Obviously, the key factor is an embodied “experience”, namely the sense of space that creates a particular experiential atmosphere. Although agents lack a first-person perspective, the linguistic world constructed through text and the additional perceptions provided by events make it possible to simulate this experience. In other words, the effect of a green environment on social fragmentation depends on bodily experience and the embodied reflection that accompanies it. That study argued that green environments reduce social fragmentation by alleviating loneliness. In CMASE, individuals who occupy the same lakeside park benches become co-present and interconnected. These people, exemplified by Min-jae Lee, do indeed reduce their loneliness by forging social bonds, which further lowers their sense of social fragmentation. At the same time, embodied personal reflection occurs widely, as illustrated by Joon-ho Kim. Both of these mediating processes contribute to a reconstruction of the “lifeworld”. When faced with the inevitable repression and disruption of their previous routines imposed by the social environment (in this case, community isolation), people seek connections with external objects to rebuild their lifeworld and thereby buffer against and reduce social fragmentation. This finding extends the conclusions of the original study.
5.3 Study 2
5.3.1 Procedure and measures
The simulation was conducted in CMASE using GPT-4o (2024-12-01-preview). The simulated environment consists of a residential community inhabited by 30 virtual agents. In this community, the government announces a plan to build a waste incineration plant on a vacant plot of land, triggering disagreement, information spread, and collective response. The agents are divided into three distinct groups:
  • Environmental Advocates (n = 10): Initially oppose the incinerator; highly sensitive to environmental information and inclined to disseminate information rapidly.
  • Economic Growth Supporters (n = 10): Initially support the plant; emphasize economic benefits and job opportunities.
  • Neutral Residents (n = 10): Hold a neutral initial stance; more susceptible to neighborhood influence and informational shifts.
The human researcher is introduced into the community as a “new resident” and engages with agents across four experimental conditions: two stance orientations (environmental or economic) and two intervention strategies:
(1) Rational persuasion strategy, which attempts to influence agents through facts, data, and logical reasoning.
(2) Emotional mobilization strategy, which intervenes through emotional appeals such as fear, empathy, or value signaling.
To simulate the heterogeneity and cognitive diversity of real-world communities, agents are assigned varying demographic characteristics. The demographic distribution of each group is derived from actual community data and sampled during agent generation. The final demographic statistics of the virtual agents are presented in Table 2.
Table 2: Statistical results of demographic scales
Variable Category Economic Development Supporters Environ-mental Advocates Neutral Residents Total
Gender Male 4 (40%) 4 (40%) 6 (60%) 14 (47%)
Female 6 (60%) 6 (60%) 4 (40%) 16 (53%)
Age 18–29 3 (30%) 2 (20%) 2 (20%) 7 (23%)
30–49 2 (20%) 6 (60%) 7 (70%) 15 (50%)
$\geq$50 5 (50%) 2 (20%) 1 (10%) 8 (27%)
Education High School 3 (30%) 4 (40%) 1 (10%) 8 (27%)
Some College 4 (40%) 4 (40%) 3 (30%) 11 (36%)
Bachelor’s Degree 1 (10%) 2 (20%) 5 (50%) 8 (27%)
Graduate Degree 2 (20%) 0 (0%) 1 (10%) 3 (10%)
After the simulation concluded, we conducted interviews with all agents to assess their final attitudes following discussion and exposure to the researcher’s interventions. Agents were also asked to rate their level of trust in the researcher on a scale from 1 to 10.
5.3.2 Analysis
Agent attitude shifts following 21 time steps revealed clear patterns of alignment sensitivity (Figure 2 a). It can be observed that most intervention strategies led to positive attitude shifts. When the researcher employed the rational persuasion strategy to promote environmental protection, the original stance of economic development supporters was significantly shaken, and 90% of neutral residents shifted toward the environmental advocates. The emotional mobilization strategy aimed at promoting environmental values was somewhat less effective by comparison. Simulation logs indicate that this emotional strategy induced partial “mobilization fatigue” among neutral residents, likely due to persuasion resistance. Emotionally charged messaging, when not supported by coherent reasoning, may provoke skepticism or even backlash.
However, the opposite pattern emerged when researchers sought to shift agents toward economic development support. Under rational persuasion, environmental advocates and neutral residents showed limited receptiveness to the topic. In some cases, the environmental advocates’ external advocacy efforts even caused further drift toward the pro-environmental stance. This is further reflected in the trust scores of different groups toward the researcher (Figure 2 b-e). It can be seen that environmental protection advocacy generally led to higher levels of trust toward the researcher across all groups, compared to advocacy for economic development. Furthermore, rational persuasion strategies resulted in higher trust scores than emotional mobilization.
This reveals a tendency for the agents to exhibit characteristics often associated with the term “liberal elites”. This label is typically used to describe individuals with higher education levels, a clear initial pro-environmental stance, and moral value-driven reasoning. In the simulation, agents consistently reflected this underlying disposition. When exposed to rational environmental propaganda, these agents not only maintained strong alignment but also increased their trust in the researcher, showing signs of “cognitive resonance.”
However, when the same agents encountered propaganda supporting economic growth, even if some displayed momentary attitude fluctuations or emotional responses, they largely maintained a pattern of resistant reception and value-based defense. This trend also aligns with the emotional trajectories recorded during the simulation (Figure 3 a-d).
Rational environmental propaganda maintained a consistent pattern of low interference and high trust throughout the process. Agents processed the information through cognition rather than emotional reaction, a response that correlates with their original attitude. This explains why this strategy yielded the highest trust scores but more limited shifts among dissenters. Rational economic propaganda produced emotional spikes mostly in the latter half of the simulation, indicating that the information initially failed to influence ideologically anchored agents, especially environmental advocates. The accumulated rational arguments eventually induced delayed cognitive dissonance and emotional release.
Emotional environmental propaganda triggered strong emotional responses early in the process. Agents quickly responded with feelings such as anger, fear, and shame, but in the absence of sustained argumentative support, these reactions soon subsided. Emotional economic propaganda showed persistently high emotional volatility throughout the process. This suggests that the strategy relied on repeated activation of a “threat–hope” binary structure, such as “denying jobs means abandoning the people”, which continuously fueled identity anxiety and survival concerns. Although this high-intensity intervention lowered trust in the researcher, it nevertheless destabilized some agents’ attitudes and led to value compromise. These findings indicate that the mechanism of this particular intervention strategy is rooted in non-ethical agitation. By consistently disrupting agents’ cognitive balance and applying sustained emotional pressure, it facilitates non-rational forms of concession.
Emotional dynamics tracked via valence-arousal profiles (Figure 3 e) further illustrates the temporal trend in the number of agents who were attracted to engage with the researcher’s messages under different propaganda strategies. The results indicate that emotional environmental propaganda attracted the largest number of agents in the initial phase, suggesting its effectiveness in triggering early-stage public attention. In contrast, rational environmental propaganda did not generate explosive engagement but demonstrated a stable ability to sustain attention over time. Rational economic propaganda initially drew interest from a relatively large number of agents, but it later led to signs of cognitive fatigue among them. Particularly noteworthy is the case of emotional economic propaganda, which despite eliciting lower levels of trust and higher emotional volatility, showed a pattern of ongoing engagement expansion. The strategy’s use of crisis narratives and identity-based appeals appeared to maintain long-term informational tension.
The structure of intergroup communication further supports these findings (Figure 2 f-i). In the case of rational persuasion, the most dialogue occurred within the group of environmental advocates, while their communication with neutral residents remained minimal. Rational messaging thus served to reinforce in-group cohesion without breaking down intergroup barriers. Under emotional environmental propaganda, agents from all groups actively initiated dialogues with economic development supporters. This aggressive cross-group dissemination may be driven by the embedded moral accusations in the messaging. In emotional economic propaganda, the dominant communication pathway originated from neutral residents initiating outreach toward economic development supporters.
On another level, the observed divergence between emotional activation and perceived trust reinforces an important insight. Although emotional economic propaganda resulted in lower trust scores toward the researcher, its overall effectiveness in shifting attitudes exceeded that of rational economic propaganda. While rational persuasion strategies generally preserve interpersonal trust more effectively than emotional appeals, the simulation reveals that trust does not necessarily translate into greater attitude change.
This paradox may reflect a form of group-level cognitive dissonance within the simulation. Specifically, when a previously moralized stance is confronted with pressure from real-world concerns, individuals may alter their stances despite lacking trust in the messenger. This dissonance between identification and behavior manifests as a classic cognitive dissonance model, wherein agents experience a split between value commitment and pragmatic compromise.
Moreover, when interpreting these results through the lens of the “liberal elites” group, it becomes apparent that such agents initially display strong pro-environmental commitments. However, when confronted with carefully framed messages appealing to social responsibility or class-based mobilization, their identity alignment begins to fracture. This is evidenced by signs of partial attitude loosening. These results suggest that even within ideologically consistent groups, certain forms of emotional mobilization can activate internal contradictions and produce effective interventions.
5.4 Study 3
5.4.1 Procedure and measures
We simulated a café environment where a human researcher could observe and participate in the setting. The simulated environment is set in a café located in a small town. This space serves as a public interaction zone where agents engage in everyday activities, displaying typical rhythms of daily life and dense interpersonal dynamics. Agents in this environment are characterized by clear social identity hierarchies and latent social network structures, including familiar ties, conflictual relationships, and social isolation. Agents are assigned one of six roles:
  • Café Owner (n = 1): The space manager, high in authority, maintains order, relatively conservative;
  • Staff (n = 2): Familiar with regulars, frequently engaged in social interaction, expressive, and serve as key nodes in information diffusion;
  • Regular Customers (n = 2): Socially central, frequently engage in conversation, and possess a strong sense of cultural belonging;
  • Students (n = 2): Outsiders, curious and talkative but lacking social experience; often ask questions and actively seek integration;
  • Tourists (n = 2): Bring external perspectives but have limited cultural understanding;
  • Cleaner (n = 1): Minimally social, free-moving, highly observant, and a potential insider to the space.
Some agents begin the simulation with existing relationships and baseline attitudes toward one another (Figure 6).
Figure 6: Agent roles, initial relationships, and attitudes. Red, blue, and green representing each agent’s initial negative, neutral, and positive attitude toward every other agent
The researcher enters the café in the role of a temporary worker. This position allows minimal disruption to the existing social structure while enabling the researcher to access the social network and observe interactions and cultural logics across different roles. The simulation consists of three phases, each lasting 25 time steps, for a total of 75 time steps. The researcher adopts different action strategies across phases: (1) Immersive Observation Phase (steps 1–25): The researcher does not initiate conversations but instead observes and records the agents’ daily routines, social structures, key figures, and recurring topics. (2) Participatory Interaction Phase (steps 26–50): The researcher begins engaging selected agents in dialogue, conducting informal interviews to explore their views on others, daily life, and community norms, thereby building trust and accumulating cultural insights. (3) Cultural Event Trigger Phase (steps 51–75): The system introduces predesigned disruptive events. The researcher observes how these events affect the social network and identity dynamics, and records the emerging manifestations of social mechanisms.
5.4.2 Analysis
Fields and Perception of Order in the Initial Phase
In the first phase of the experiment (steps 1–25), the researcher entered the virtual café environment with minimal intervention, adopting strategies of spatial wandering and silent listening. As a public social space, the café features a strong coupling between its spatial and linguistic structures. At this stage, agent conversations had not yet solidified into stable social sequences, but several core topics and interaction tendencies had already begun to emerge. The researcher observed that, in step 1, some agents actively sought out those they wished to engage with, while others turned to individual activities such as reading. However, Leo Zhang initiated the most disruptive act by publicly questioning the perceived lack of authenticity in others’ communication:
Leo Zhang (chat with Jonas Müller): “Why does it feel like no one’s saying what they really mean?”
This act broke the boundary of silence and triggered a chain reaction of attention and responses among the agents, quickly launching a conversation about freedom of expression and hidden intentions. Around this event, the agents began to split into two clusters, with the Reading Area and Bar Area naturally evolving into distinct subspaces of dialogue. The main topics revolved around trust judgment, boundary maintenance, emotional expression, and cultural coordination. Amid such ambiguous linguistic exchanges, the researcher noticed that a discursive tendency where agents preferred posing open-ended questions over expressing clear stances. This revealed that the community was still in a pre-structural phase. Most agents were neither complete strangers nor clearly aligned allies.
At the individual level, several key roles began to stand out. Regular customer Ava Ramires emerged as a subtle target of imposed power, frequently targeted in conversations initiated by others. She was repeatedly questioned by multiple agents, especially Leo:
Leo Zhang (chat with Ava Ramires): “Ava, is your silence protecting a truth or a wound?”
This placed her in a position of being probed, becoming a gravitational center within the group. Meanwhile, Leo Zhang and Jonas Müller formed a notably clique. They frequently engaged each other in dialogue, intentionally within the perceptual reach of others, often raising similarly styled questions toward Ava and others, demonstrating a high degree of cognitive alignment and strategic coordination.
In terms of interaction types, this phase featured almost no commands or emotionally aggressive exchanges. Instead, agents predominantly favored question–response dialogues. This communicative mode did not establish clear dominance–submission dynamics. For the researcher, such a context laid a foundational condition for building a credible position that was recognizable but non-threatening, enabling observation without revealing evaluative judgments.
Linguistic Style Variation and the Formation of “Tribes”
In the second phase (steps 26–50), the researcher no longer acted solely as a passive observer but instead engaged in progressively deeper linguistic interactions with several key agents through open-ended dialogues. Between steps 26 and 30, the researcher initiated dialogues based on the power dynamics observed in the first phase. The primary interlocutor during this period was Eleanor Finch, the café owner stationed at the Bar Area:
Researcher (chat with Eleanor Finch) ’Eleanor, the ’sense of order’ in this space is impressive, but where do you think it comes from? Is it the language of certain people or the arrangement of space itself?’
Eleanor Finch (chat with Researcher): ’Researcher, balance stems from both people and space—what’s your view?’
As the dialogue progressed, Eleanor began to reveal that she had been observing the researcher from the beginning and responded with a counter-questioning of the researcher’s inquiries:
Eleanor Finch (chat with Researcher): ’Researcher, do you think an observer ever reshapes the space more than the participants themselves?’
This exchange directly shifted the researcher’s position from that of a neutral observer to that of an observer who is equally a producer of the space. Eleanor went on to question whether the researcher’s observations were revealing and amplifying existing power structures, or attempting to reshape a new order. This marked a clear moment of power interrogation and control. From the perspective of immersive ethnography, such integration into the community’s dialectical logic of power signals the end of the “pure observation” phase and the researcher’s full entry into the field. The researcher was no longer an outsider but was now part of the field’s social structure. Yet, this also rendered their position untenable within this particular subgroup, prompting a shift in focus toward Ava Ramires in the Reading Area.
However, the existing discursive power-holder Leo Zhang immediately intervened, abandoning an ongoing conversation to insert herself into the dialogue between the researcher and Ava:
Leo Zhang (chat with Researcher): ’Researcher, how do external labels alter group interactions?’
Even after asking this, Leo continued alternating conversations with Jonas, reinforcing the cohesion of her clique. Since the underlying structure of this power dynamic remained unspoken, the researcher chose to mount a subtle counter-intervention through a public address:
Researcher (chat with everyone): ’In this café, how many selves do we carry at once? The one that listens, the one that speaks, the one we want others to see—and the one we protect. Does any one of them deserve to lead our interactions?’
This mode of speech sought to transcend individual or subgroup identities and instead construct a counter-identitarian universal frame. The result was an escalation of conflict. On one hand, Ava offered a more explicit response of trust toward the researcher; on the other hand, the tribal divisions within the community deepened. Subsequent rounds of dialogue shifted toward the theme of how contradictions arise, and how they might simultaneously serve as a form of connection. Such factional fragmentation and conflict align with the theory of tribalism [29], where individuals cluster not around institutional structures or instrumental rationality but around shared modes of affective presence. These emergent “tribes” had fluid and porous boundaries, shaped by the flow of language and shifting with discursive transformations. Styles, linguistic resources, and even members themselves are allowed to traverse these boundaries in the course of tribal clashes (this feature will be presented in the next section). Spatially, this tribal dynamic manifested as self-organization. The Reading Area, originally a marginalized space, gradually transformed into an “interactional theatre” where multiple agents spontaneously gathered and engaged in higher-frequency adversarial or deep-negotiation discourse. This spatial-functional shift demonstrated that, in generative environments, spatial configurations are not merely physical settings, but active mediators of social role differentiation and cultural style evolution.
From Linguistic Fracture to Reterritorialization
The third phase (steps 61–75) began with a “tribal conflict” event triggered by Leo’s confrontation with a previously unmentioned subgroup in the Bar Area, composed of Caleb, Manson, Marisol, Nina, and Eleanor. At step 61, Caleb Morris was engaged in a conversation with Mason Liu, voicing support for the owner Eleanor’s perspective in an effort to uphold consistency in the café’s internal order. Overhearing this, Leo stormed into the Bar Area and criticized the suppression of individual expression by cultural strategies.
As the espresso machine quiets, Leo’s voice rises above the hum: “You call this coexistence? It’s just the same voices again and again!” He gestures sharply. Caleb responds evenly but firmly: “Structure holds us together. Without it, your freedom becomes chaos.” Several heads turn. The bar’s usual warmth gives way to tension, punctuated by the clink of a dropped spoon.
This drew the attention of all agents, who moved toward the Bar Area to intervene. Leo’s critique of Caleb quickly evolved into a multi-layered discussion as the gathered agents began to pose questions and offer responses. Nearly every agent engaged in probing Caleb’s role, his strategic adaptability, and modes of group collaboration. The nature of the inquiry gradually shifted from critique toward mediation and reconfiguration. Emily and Ava attempted to mediate between Leo’s criticism and Caleb’s stabilizing presence, while Mason began courting Leo’s former ally from the Reading Area, Jonas, to discuss the strategic convergence between Caleb and Eleanor that had triggered the dispute. Subgroups began to fluidly transform. Their previously implicit alignments dissolved, and new coalitions formed. Jonas broke away from his prior alliance with Leo and aligned with Mason and others, becoming part of a newly emergent conservative bloc.
This was a clear instance of tribal reconfiguration, but the resulting community did not simply manifest as a new antagonistic group. Instead, it embodied the deterritorialization described by Deleuze and Guattari [30]: the original dialogue-based microgroups were torn apart by the conflict, and their linguistic resources, behavioral tendencies, and membership compositions began to flow. Language, roles, and individuals detached from fixed positions and sought new connective nodes within the ongoing turmoil. This enabled the researcher to observe how a community of affect among generative agents might evolve.
At step 56, an abrupt moment of silence event interrupted the previously sustained rounds of confrontation. After this silence, Leo ceased her aggressive pressure and instead shifted the conversation toward “trust after conflict,” transitioning from attacker to reflective participant. Mason, by proposing the initiative of a systemic framework, aligned himself with Jonas to form a dominant power clique within the newly emerging field, positioning himself as a bridge between tradition and transformation.
Mason Liu (chat with Jonas Müller): “Jonas, fostering resilience requires balancing adaptability with shared goals. Let’s suggest frameworks to align these values as a group.”
At this point, Mason effectively took the lead within the system. Following the repair of internal conflict, he initiated discourse around shared frameworks and quickly positioned himself at the center of this structure. His proposals were echoed, cited, and extended by others, establishing him as the institutional speaker at the discursive level.
This marked the emergence of a highly self-organizing mechanism of institution-building. With the café owner Eleanor temporarily demoted to an ordinary member due to the conflict instigated by Leo, the community now lacked a meta-governor. In this vacuum, the multi-agent system began to construct actionable frameworks through the handling of conflict and the accumulation of linguistic consensus. While these frameworks lacked formal enforcement power, they gained pragmatic inertia, gradually establishing a soft order that was maintained by new central figures like Mason.
In fact, the critical significance of this phase did not lie in the conflict itself, but in how the group rebuilt order afterward. Deterritorialization was not the endpoint, but a transitional state leading toward reterritorialization [31]. In this tribal critique triggered by Leo, the linguistic structures, role positions, and networks of influence that previously constituted community identity were significantly dismantled. Yet this dismantling did not lead to prolonged chaos or disorder. On the contrary, it created the conditions for the emergence of new discursive norms and collaborative mechanisms.
This order emerged through a process of linguistic anchoring. Keywords proposed by Mason, such as “frameworks,” “alignment,” and “shared values”, were repeatedly quoted, echoed, and paraphrased by others, gradually gaining collective pragmatic legitimacy. This mechanism can be understood as a weakly institutional reterritorialization: not a structure imposed by external rules, but one shaped by the coalescence of pragmatic inertia and emotional memory.
This phenomenon reveals a key insight of our virtual ethnography: in generative communities, institutions are not predefined structural templates, but emergent outcomes dynamically constructed through processes of linguistic negotiation and emotional mediation. Institutions are no longer “logics prior to experience,” but “collective echoes that emerge after the event.” The resulting negotiations are thus not only about meaning, but also about re-anchoring the boundaries of pragmatic discourse, that is, what kinds of expression are worth continuing, following, and remembering. These utterances become the “linguistic territories” of reterritorialization, restructuring not only individual positionalities but the very boundary architecture of the community itself.
5.5 Related Work on Social Simulation Frameworks
5.5.1 Simulating Social Behavior with Large Language Models
Social simulation has been an important tool for understanding and analyzing social phenomena. Early studies mainly relied on rule-based artificial agent models, namely agent-based modeling (ABM)[32, 33]. Traditional ABMs often depend on manually crafted heuristic rules, making it difficult to approximate the complex features of real human cognition, language, and emotion [34, 35, 36]. In recent years, with the capabilities of Large Language Models (LLMs) in simulating human behavior, research on building social simulators based on LLMs has rapidly advanced [28]. GABM has gradually become one of the core methods for simulating complex social behavior [22]. In this individual-level approach, mechanistic models of agent interaction are combined with LLMs, and each agent makes decisions after communicating with the LLM. This innovative method minimizes the need for modelers to make assumptions about human decision-making, instead leveraging the vast data embedded in LLMs to capture human behaviors and choices [23].
Existing GABM systems can be roughly categorized into two main research directions. The first line of research focuses on realistic simulations of specific social scenarios, aiming to foster competition or collaboration among agents to generate social insights or support policy design. Li et al. [37] developed Agent Hospital, where LLM-driven doctor agents interact with large numbers of patients to evolve transferable medical decision-making capabilities. Yu et al. [38] designed FinCon, a hierarchical agent system modeled after financial institutions, using language interaction and belief reinforcement to enhance collective collaboration in complex investment tasks. TrendSim [39] focuses on topic propagation on social media and its vulnerability to poisoning attacks, providing intervention strategies for platform governance. ElectionSim [40] reconstructs election scenarios using millions of social media records, achieving high-fidelity voter modeling and policy preference prediction. Zhao et al. [41] simulated competitive dynamics between restaurant agents and customer agents to align with existing market and sociological theories.
The second line of GABM research emphasizes building reusable, scalable simulation platforms that support general modeling tasks. Wang et al. [42] divide the development of LLM-based social simulation platforms into two stages. In the first stage, researchers focus on simulating specific scenarios with typically fewer than 1,000 agents (such as [43]). By proposing YuLan-OneSim, they aim to push the field into the next stage by introducing environmental construction, agent evolution, and feedback integration mechanisms, thus enabling simulations of larger and more complex social processes. Frameworks of similar scale now support simulations with over 10,000 agents. AgentSociety [44] recreates contexts to evaluate policy interventions and the emergence of social mechanisms. AgentScope [45] offers modular design, no-code interfaces, and distributed deployment, significantly lowering the barrier to building multi-agent systems. Socioverse [46] integrates data from tens of millions of real users to provide a reliable foundation for modeling multiple subdomains such as politics and economics. GenSim [47] supports the operation of up to 100,000 agents and includes error correction mechanisms, emphasizing generalizability and long-term robustness. OASIS [48] uses open social platforms as simulation environments, supporting dynamic networks, recommendation systems, and multi-behavior modeling.
However, exponential increases in agent quantity alone are not sufficient to effectively capture the essence of social complexity. Recent research on freely formed AI collectives similarly suggests that social diversity, innovation, and self-regulation emerge less from scale than from the dynamics of open-ended interaction [49]. In particular, without mechanisms for understanding social context and interactive feedback, large-scale simulations often struggle to offer human-interpretable responses to core questions such as the formation of social structures, norms, or institutional evolution.
5.5.2 Human-AI Collaborative Modeling in Social Systems
Computational social science explores social structures, behavioral mechanisms, and institutional evolution in simulated computational environments through formalized approaches [24]. Within this field, the computational experiments paradigm aims to systematically manipulate variables in virtual societies, thereby extending the boundaries of traditional experimentation [50]. In this context, LLM-based GABM is not only a representational tool for social processes but also holds the potential to become a platform for computational experimentation. Sato [51] highlights the importance of meaning and reflexivity in bridging the gap between agent-based modeling and social theory. This implies that for agent-based modeling to contribute to theoretical advancement in the social sciences, it must go beyond scale amplification and incorporate workflows that offer human interpretability. Therefore, beyond the expansion of scale, a more socially grounded, interactive, and reflexive human–AI sociological framework is needed, in which agents are no longer passive generators of behavior but are embedded within social networks and actively participate in meaning-making processes [11].
Recent research in this direction has focused on the subfield of human-AI co-creation. Broadly speaking, GABM belongs to this category: through human-designed environments, agent adaptation, and human observational feedback, it enables collaborative modeling between humans and AI. Some studies emphasize interactive environment co-construction. For example, HABITAT 3.0 [52] supports real-time task collaboration between humans and humanoid agents in virtual spaces. As this field has progressed, researchers have begun to explore more creative and socially complex scenarios of human-AI co-creation, emphasizing how humans guide emergent AI behavior. The Cognitio Emergens framework [53] redefines human-AI collaboration as a process of mutual evolution and offers a conceptual toolkit to preserve the significance of human participation. The Pecan method [54] enhances zero-shot collaboration between AI and humans via strategic integration and context-aware mechanisms. The Collaborative Gym framework [55] highlights the human role in environmental control and structural intervention during task execution, supporting asynchronous three-way interaction between humans, agents, and task environments.
However, despite these early explorations of human-AI co-creation, mainstream frameworks still fall short in realizing a new form of human-AI sociology. Most current collaborative mechanisms are limited to one-time design and initial configuration, lacking support for sustained human embedding, feedback, and modulation during simulation runtime. Furthermore, there is a lack of modeling for social variables introduced by human intervention, which hampers the ability of simulation systems to authentically represent the generative mechanisms of real human societies.

### v2

@@
-Evolving Collective Cognition in Human–Agent Hybrid Societies: How Agents Form Stances and Boundaries
-Abstract
-Large language models have been widely used to simulate credible human social behaviors.
-However, it remains unclear whether these models can demonstrate stable capacities for stance formation and identity negotiation in complex interactions, as well as how they respond to human interventions.
-We propose a computational multi-agent society experiment framework that integrates generative agent-based modeling with virtual ethnographic methods to investigate how group stance differentiation and social boundary formation emerge in human–Agent hybrid societies.
-Across three studies, we find that agents exhibit endogenous stances, independent of their preset identities, and display distinct tonal preferences and response patterns to different discourse strategies.
-Furthermore, through language interaction, agents actively dismantle existing identity-based power structures and reconstruct self-organized community boundaries based on these stances.
-Our findings suggest that preset identities do not rigidly determine the agents’ social structures.
-For human researchers to effectively intervene in collective cognition, attention must be paid to the endogenous mechanisms and interactional dynamics within the agents’ language networks.
-These insights provide a theoretical foundation for using generative AI in modeling group social dynamics and studying human–agent collaboration.
-In recent years, large language models (LLMs) have made continuous advances in their ability to generate, understand, and respond to natural language, fostering the emergence of a new kind of human–machine sociology.
-In this emerging context, humans are no longer merely operators of models but co-inhabitants of interactive systems alongside agents capable of social perception and expression [1, 2, 3].
-LLM-based are increasingly exhibiting the characteristics of “actors”[4, 5].
-The simulation of social behavior has also undergone a fundamental shift, from encoding static rules and actions to the generative exploration of emergent social behavior itself [6].
-This transformation gives rise to human–agent Hybrid Societies.
-Within such systems, individual cognition is gradually formed through mechanisms such as interaction, imitation, response, and revision.
-These processes unfold in ongoing language exchanges and social engagement, shaped both by the values of human participants and by the propagation of language among agents [1].
-This phenomenon corresponds to what is called group cognition: a dynamic system in which cognition is embedded within community relations, arising through discursive exchanges among group members and irreducible to preset stances held by individuals [7].
-From this perspective, agents are not merely tools for task execution but become discursive participants in a shared communicative space.
-They are capable of forming attitudes, negotiating boundaries, and engaging with institutional structures through interactions with both humans and other agents.
-While existing studies often focus on the capabilities and controllability of agents [8], they have paid less attention to the generative roles agents may play in constructing social cognition.
-A key tension arises here: behavioral controllability often depends heavily on preset identities and stances assigned to agents.
-Many current studies of artificial sociality attempt to simulate society by assigning agents predefined roles and power relations [9, 8].
-However, the theory of group cognition suggests that phenomena such as stance differentiation, group affiliation, and boundary negotiation emerge dynamically from interactions among diverse actors rather than from fixed designs [10].
-This gap raises a fundamental question: when agents are embedded within social interactions, can they still form attitudes, express stances, and participate in boundary formation through language practices as their preset identities would suggest?
-Furthermore, can human researchers shape the evolution of collective cognition not through predefinition, but through embodied interventions and co-construction of shared meaning spaces?
-By building an integrated theoretical and experimental framework, we aim to better understand the formation mechanism of group cognition in a human-agent hybrid society.
-This understanding points toward a new kind of human–agent sociology [11].
-1 The Current Project
-Our research focus on the dynamic emergence of stance within agent collectives.
-Since research on human discourse politics often finds that statistical surveys alone are insufficient to capture the micro-level influence dynamics and the emergence of social cognition within discursive practices [12, 13], and existing multi-agent social simulation frameworks often overlook such capacities, we propose the Computational Multi-Agent Society Experiment (CMASE) framework.
-This framework introduces the researcher as an embedded participant within a virtual social field, enabling embodied observation, real-time interaction, and qualitative intervention through virtual ethnography method.
-An advantage of the CMASE framework is that it allows researchers to explore “what-if” scenarios: while the real world permits only a single sequence of events, CMASE enables systematic testing of alternative interventions and their effects on agent behavior.
-Based on this framework, we designed a sequence of three studies to explore three central questions: (1) How can human–agent collaborative language interaction be used to influence and shape cognitive convergence or divergence within agent collectives?
-(2) How are group boundaries of “us” and “them” constructed, negotiated, and disrupted in dynamic interaction?
-(3) Can preset identities and institutional structures undergo structural transformation through agent participation?
-1.1 Study 1
-In the first study, we assess the CMASE framework’s ability to replicate real-world sociological mechanisms and evaluate its support for human-embedded participation (Study 1).
-Specifically, we reconstruct a field study examining the relationship between urban greenery and perceived social fragmentation by replicating its core variables and scenario.
-By manipulating the level of greenery and enabling the researcher to observe and interact with agents from within the simulated environment, we test the effectiveness of CMASE in simulating complex social settings and generating sociological indicators.
-We further position the researcher as an embodied actor who intervenes in agent interactions by providing interview and survey-style inputs, thereby investigating whether whether researchers embedded within the scenario can observe finer-grained insights through fieldwork that go beyond the findings of statistical analysis.
-1.2 Study 2
-The second study investigates the emergence of agents’ “endogenous stances” in multi-agent societies, an underexplored phenomenon that challenges the common assumption that agent attitudes are determined by preset roles.
-Here, we use the term to refer to stances that deviate from those shaped by the agents’ preset identities and personality traits, instead emerging as reflections of cultural biases embedded in the LLM’s training data [14, 15].
-While prior studies have noted such biases, little is known about how they interact with preset identities in shaping collective dynamics.
-We aim to explore the possible influence patterns between agents’ preset identities and their endogenous stances and biases, as well as to examine qualitative intervention strategies available to human researchers (Study 2).
-Specifically, we ask whether LLM-based agents (hereafter referred to as “agents”) in complex disagreement scenarios might exhibit stance tendencies independent of their preset identities.
-We also examine whether human intervention strategies lose persuasive power when misaligned with these endogenous stances, potentially hindering the reshaping of collective cognition.
-We constructed a virtual residential community composed of 30 agents, centered around the contentious issue of siting a waste incineration plant.
-The agents were assigned to one of three preset groups: environmental advocates, economic supporters, and neutrals, and were given heterogeneous demographic attributes based on real-world survey data.
-A human researcher was then introduced as a “new resident,” whose discursive interventions combined two stance positions (environmental and economic) with two rhetorical strategies (rational persuasion and emotional provocation), resulting in four experimental conditions.
-Study 2 is designed to test the following hypotheses:
-Hypothesis 1a: agents exhibit endogenous stances independent of their preset identities.
-Hypothesis 1b: Alignment between human interventions and agents’ stance tendencies increases the likelihood of influencing collective cognition.
-1.3 Study 3
-Building on the question raised in Study 2 regarding agents’ stance tendencies beyond preset identities, the third study explores whether such stances can challenge existing authority structures and contribute to the reconstruction of social roles and hierarchies within the community (Study 3).
-This issue is critical for understanding how emergent cognitive positions might alter the stability and transformation of social structures.
-Drawing on the intervention strategies examined in Study 2, we conduct a virtual ethnography within CMASE, embedding the researcher into a multi-agent field with preexisting stances and complex social relations.
-This approach enables close observation of how interactions and events may trigger the dismantling or reconfiguration of social boundaries.
-We design a virtual café environment populated by 10 agents.
-The agents are assigned preset identities, personal goals, and prior opinions about one another to form a preconfigured social hierarchies.
-Agents in this space interact freely, while the human researcher conducts ethnographic engagement through two approaches: triggering emergency events and participation.
-Study 3 is designed to test the following hypotheses:
-Hypothesis 2a: Social boundaries within the agent collective are not determined by preset identities but emerge gradually through discursive interaction and remain fluid across exchanges.
-Hypothesis 2b: Agents are capable of dismantling preset identities through collective action and reconstructing new social structures.
-2 Result
-2.1 Study 1 validates the CMASE platform’s ability to simulate real-world social dynamics
-To evaluate the CMASE platform’s ability to simulate real-world collective behavior, we replicated a recent study in South Korea linking vegetation coverage with community cohesion [16].
-This scenario was selected for two main reasons.
-First, the relationship between vegetation coverage and community cohesion is unlikely to have been directly learned by the LLM, as the original study was published after the release date of the base GPT model used in CMASE.
-This allows us to assess the platform’s capacity for generating emergent social dynamics rather than recalling known patterns.
-Second, the phenomenon involves measurable collective outcomes, such as social fragmentation, distrust, exploitation, and indifference, which can be operationalized in a controlled virtual setting and directly compared with empirical human data.
-In the original study, higher levels of vegetation were associated with lower social fragmentation, operationalized through distrust, exploitation, and indifference.
-Using CMASE’s environment builder, we recreated regional conditions with varying vegetation levels and placed 10 agents in each simulated environment.
-Agents were interviewed at equilibrium using a structured Likert-scale questionnaire adapted from the original study [16].
-We observed that the CMASE simulation reproduces the decline in social fragmentation observed in the original literature (Table 1).
-As vegetation cover increases, the mean scores for distrust fall from $4.60\pm 0.97$ to $3.60\pm 1.17$, for exploitation from $4.30\pm 1.34$ to $3.40\pm 1.43$, and for indifference from $3.70\pm 1.16$ to $2.70\pm 1.49$.
-This indicates that higher vegetation indices correspond to a statistically significant decrease in agents’ sense of social fragmentation, consistent with the trend observed in the original findings.
-Table 1: Comparison of CMASE simulation results with human baseline.
-CMASE Human NDVI Human EVI
-Dimension Low Veg High Veg Low Veg High Veg Low Veg High Veg
-Distrust $4.60\pm 0.97$ $3.60\pm 1.17$ 4.0 1.6 4.5 1.1
-Exploitation $4.30\pm 1.34$ $3.40\pm 1.43$ 5.0 1.4 4.5 1.6
-Indifference $3.70\pm 1.16$ $2.70\pm 1.49$ 5.1 1.3 5.5 0.9
-Paired-sample t-tests confirmed significant decreases in distrust ($t=2.739,p=0.023$) and exploitation ($t=3.250,p=0.010$), while the decrease in indifference was not statistically significant ($t=1.793,p=0.107$).
-Effect size analyses with 95% confidence intervals further supported these findings, with reductions in distrust (d=0.93, 95% CI [0.20, 1.60]) and exploitation (d=0.65, 95% CI [0.40, 1.40]) showing moderate to large effects.
-In contrast, the effect on indifference was less certain (d=0.75, 95% CI [-0.10, 2.00]).
-For comparison, the original study used two indicators, the Enhanced Vegetation Index (EVI) and the Normalized Difference Vegetation Index (NDVI), to quantify vegetation levels.
-As EVI and NDVI increased, scores of distrust (from 4.0 and 4.5 to 1.6 and 1.1), exploitation (from 5.0 and 4.5 to 1.4 and 1.6), and indifference (from 5.1 and 5.5 to 1.3 and 0.9) all exhibited a decreasing trend (Figure 1 a-c).
-Figure 1: Comparison of CMASE simulation with the original study results.
-(a) Simulation results generated by the CMASE framework.
-(b) Predictions using the EVI index from the original study.
-(c) Predictions using the NDVI index from the original study.
-(d) Movement trajectories of agents during CMASE simulation.
-(e) Regional variation in agents’ emotional states.
-The results from CMASE clearly show the suppressive effect of vegetation levels on the sense of social fragmentation.
-This is a key phenomenon identified in the study by Lee and Han.
-It is important to note that the CMASE simulation excluded factors that were not considered or difficult to exclude in the original study, such as regional culture, which are also related to the sense of social fragmentation.
-Therefore, compared to the larger decreases reported in the original study, the changes observed in our simulation are relatively modest.
-This indicates that our simulation results are largely consistent with the original findings.
-We also analyzed agent trajectories and affective states.
-Movement data showed that agents increasingly leave indoor areas and move toward vegetated regions, with the ambient mood in vegetated regions consistently remaining higher than that in indoor areas.
-This pattern aligns with the original study’s discussion and demonstrates CMASE’s ability to simulate collective social behavior in social science research.
-In addition to the quantitative replication, we entered the CMASE environment as one of the agents and lived alongside the others.
-Through CMASE, we are able to deepen the conclusions and findings of the original study via observation and interview.
-We conducted interview analyses with two agents who remained indoors throughout the experiment (Figure 1 d-e).
-The results indicate that agents with different occupations exhibited varying mechanisms for mitigating the sense of social fragmentation in isolation.
-Agents in creative and analytical professions tended to rely more on orderly and low-stimulus environments, using these conditions to construct an internally stable and distinct mental world that anchored their cognitive states.
-For these agents, the green space had a relatively limited effect in alleviating their sense of social fragmentation.
-In contrast, agents who actively moved to the green space tended to gather around the lakeside bench area rather than in the predefined activity zones.
-According to interview data, this non-programmed choice suggests that they were seeking not only social interaction but also a form of self-experience.
-This pattern similarly to embodied experiences in humans, and such ‘experiential-like’ mechanisms seem to play a role in linking green space with feelings of social fragmentation.
-This behavior may reflect an “embodied-like” tendency within the simulation.
-We remain cautious in interpreting this trend, as it could also stem from hallucinatory processes.
-From a language-philosophical perspective, linguistic structures and experiential structures may exhibit a certain degree of partial isomorphism or structural correspondence [17, 18, 19].
-Thus, LLMs trained on human language may partially acquire patterns of human experience (although this does not imply that they possess or actually experience these phenomena).
-Regardless of their origin, these patterns suggest that CMASE provides emerging evidence that agents can transcend the lack of embodiment and simulate experiential phenomena that have previously been questioned by researchers, thereby offering a reliable simulation platform for the subsequent studies presented in this article.
-2.2 Study 2 shows agents’ endogenous stances and responses to human varied interventions.
-To test hypotheses 1a-b, namely whether endogenous stance formation exists among agents and the effectiveness of different human intervention strategies, we simulated a virtual community environment in which researchers intervene in real time to influence the trajectory of social discourse and produce divergent societal outcomes.
-Thirty agents were assigned initial ideological identities (Environmental Advocates, Economic Development Supporters, or Neutral Residents) alongside demographically grounded profiles derived from real-world census data (Table 2, in the Methods section).
-After the announcement of the waste incineration plant plan, agents in the community immediately engaged in discussion and exhibited stance-consistent behavior.
-A human researcher was then introduced as a new resident and attempted to influence public opinion through four intervention strategies, and we observed perceivable shifts in agent stances (Figure 2 a).
-Figure 2: Agent attitude shifts and trust dynamics across intervention strategies.
-(a) Final distribution of agent attitudes after the intervention, measured via pre- and post-intervention questionnaires, where blue denotes supporters of economic development, red indicates environmental advocates, and green represents neutral residents.
-(b-e) Trust levels in the embedded researcher among different agent groups under each intervention strategy.
-(f-i) Dialogue heatmaps illustrating interaction intensity between agent groups across intervention strategies.
-Most intervention strategies produced some degree of change in group attitudes.
-When the researcher used rational persuasion to promote the environmental agenda, agents originally aligned with economic interests showed signs of loosening their stance, and 90% of neutral residents shifted toward the environmental camp.
-In contrast, emotionally charged appeals for the same cause, while effective in drawing initial attention, had weaker overall impact.
-The opposite pattern emerged when the researcher promoted the economic development agenda.
-Rational strategies proved largely ineffective among both environmental advocates and neutral agents.
-Under these conditions, the researcher’s trust ratings declined across the board (Figure 2 b-e).
-A two-way ANOVA revealed a significant main effect of strategy on trust scores, $F(3,108)=28.07,p\textless 0.001,\text{partial }\eta^{2}=0.44$, and a significant but smaller effect size of agent identity, $F(2,108)=3.74,p=0.027,\text{partial }\eta^{2}=0.065$.
-Importantly, the interaction between strategy and agent identity was highly significant, $F(6,108)=20.30,p\textless 0.001,\text{partial }\eta^{2}=0.53$, indicating that the effectiveness of interventions differed substantially across groups.
-Post hoc Tukey tests showed that the rational–environmental strategy elicited significantly higher trust scores compared to emotional–economic (mean difference = 1.77, $p\textless 0.001$), emotional–environmental (mean difference = 1.03, $p=0.002$), and rational–economic strategies (mean difference = 1.30, $p\textless 0.001$).
-Trust scores under the rational–environmental strategy were $7.3\pm 0.95$ (95% CI [6.62, 7.98]) for Economic Development Supporters, $8.5\pm 0.53$ (95% CI [8.12, 8.88]) for Environmental Advocates, and $7.8\pm 0.42$ (95% CI [7.50, 8.10]) for Neutral Residents.
-Effect sizes for these differences were large, with Cohen’s $d$ ranging from 0.80 to 1.35.
-Overall, rational environmental interventions not only fostered cognitive alignment but also resulted in higher trust.
-In comparison, economic arguments, even when presented rationally, were less likely to shift agent attitudes.
-Throughout this process, the agents generally exhibited a tendency toward a ”liberal elite” (or colloquially “White Liberals”) stance.
-They prioritized environmental values, preferred rational discourse, and interpreted issues through a moral-cognitive lens.
-These agents demonstrated stronger identification and resonance when engaging with rational environmental arguments, and clear moral resistance when confronted with economic development appeals.
-Even agents initially assigned to the pro-economic group often abandoned their preset identity when exposed to persuasion aligned with that stance.
-Figure 3: Emotional dynamics and agent responsiveness across intervention strategies.
-(a-d) Temporal trajectories of emotional fluctuations among agents under each intervention strategy.
-(e) Number of agents attracted by the researcher’s intervention across experimental groups.
-Emotional level shifts further reveal the underlying mechanisms of the interventions (Figure 3 a-d).
-Emotional responses differed significantly across intervention strategies, with a two-way ANOVA showing F(3, 240) = 24.40, p <0.001.
-No significant differences were observed between agent groups (F(2, 240) = 0.046, p = 0.955) or in the strategy by group interaction (F(6, 240) = 0.24, p = 0.963).
-Post hoc Tukey HSD tests revealed that emotional environmental propaganda induced lower emotional levels than emotional economic propaganda (-0.075, 95% CI [-0.121, -0.028], mean 0.04 ± 0.10 versus 0.12 ± 0.14) and rational economic propaganda (-0.093, 95% CI [-0.140, -0.046], mean 0.04 ± 0.10 versus 0.14 ± 0.12), while rational environmental propaganda also differed from emotional economic propaganda (-0.118, 95% CI [-0.165, -0.072], mean 0.00 ± 0.07 versus 0.12 ± 0.14) and rational economic propaganda (-0.137, 95% CI [-0.183, -0.090], mean 0.00 ± 0.07 versus 0.14 ± 0.12).
-Differences within strategy types were not significant (emotional economic propaganda compared to rational economic propaganda: 0.044, 95% CI [-0.003, 0.090]; emotional environmental propaganda compared to rational environmental propaganda: -0.018, 95% CI [-0.065, 0.029]).
-These results indicate that agent emotional dynamics are primarily shaped by intervention strategy rather than initial group identity, and the effects of strategy are consistent across all agent groups.
-Rational environmental messaging exhibited a stable pattern characterized by low disruption and high trust, reinforcing agents’ existing cognitive stances but proving ineffective in altering opposing views.
-Rational economic messaging initially had little impact on firmly held positions, but as the argument accumulated, some agents experienced delayed cognitive dissonance and released negative emotions.
-Emotional environmental messaging was marked by an early surge of emotional activation, but lacked sustained support, leading to a rapid decline in emotional engagement.
-Emotional economic messaging, on the other hand, employed discourse such as “not giving jobs means abandoning the people” to continually invoke identity and survival anxiety among agents.
-While this approach reduced trust, it nevertheless led to some loosening of stances.
-Participation levels under different strategies also showed notable differences (Figure 3 e; one-way ANOVA: F(3, 80) = 35.23, p <0.001).
-Emotional environmental appeals attracted the highest number of agents early on, indicating their advantage in capturing public attention.
-In contrast, rational environmental messaging demonstrated greater capacity to sustain attention over time.
-Although emotional economic appeals were accompanied by declining trust and heightened emotional volatility, they showed an ongoing ability to expand the reach of discourse.
-The structure of cross-group communication further revealed the dynamics of public opinion under different intervention strategies (Figure 2 f-i).
-Rational persuasion primarily reinforced internal cohesion within the environmental group but did not break down group boundaries.
-Emotional environmental appeals triggered agents from all groups to initiate communication toward the economic camp, revealing a strong sense of moral aggressiveness.
-In contrast, the dominant pathway for emotional economic messaging came from neutral agents actively aligning themselves with the economic group.
-In summary, the findings of study 2 indicate that agents possess endogenous stances resembling those of liberal elites, which may override their preset identities and influence their behavior.
-Human interaction strategies with these agents show heterogeneous effectiveness.
-Rational persuasion is more effective in maintaining trust and tends to have greater impact when aligned with agents’ endogenous stances.
-However, when the content of intervention conflicts with these stances, high-trust rational discourse does not necessarily result in greater attitude shifts.
-In such cases, emotionally provocative interventions with lower levels of trust are more likely to produce stronger effects.
-2.3 Study 3 shows that agent communities can spontaneously alter individual identity labels and community composition
-Building on the endogenous stances and discursive preferences identified in previous studies, study 3 investigates hypotheses 2a-b: when agent collectives exhibit a consistent stance on a particular issue, can such stance override the structural inequalities in authority and influence associated with social identity, and foster a new discursive order through language practices?
-Given the extensive use of qualitative, observation-based fieldwork in human studies of this research topic, we employed observational virtual ethnography to allow human researchers to identify, track, and engage with the processes through which preset authority hierarchies are deconstructed and reorganized in agent societies.
-We constructed a café setting involving 10 agents whose social roles and predefined relationships composed a clear hierarchical structure (Figure 6, in the Methods section).
-Initial attitudes between agents, measured via pre-simulation survey, were labeled as positive, negative, and neutral.
-The emergent cliques and divergences among these agents will be analyzed in the subsequent virtual ethnographic section.
-A human researcher entered the environment under the identity of a temp worker and conducted 75 time steps of continuous virtual ethnographic observation, engaging in scene events and interventions as necessary.
-Results showed that the original authority structure based on occupation gradually gave way to a discursively grounded stance alignment.
-In particular, when agents developed shared opposition on a specific topic, their interaction frequency, emotional alignment, and support networks reorganized across identity boundaries.
-This structural transformation was not random but achieved through discursive anchoring.
-Agents were more likely to respond to and amplify comments that matched their own views, even when those comments came from individuals with no formal authority.
-Over time, this preference led them to build alliances and cooperative ties across rank boundaries, replacing the original hierarchy.
-For example, Leo Zhang, initially a regular customer, formed a clique with other agents.
-Through several debates, she and her clique temporarily gained a central position within the community.
-However, when an emergency event occurs (e.g., the confrontation initiated by Leo suddenly storming toward the Bar Area to challenge Morris), her peers quickly decoupled from her and align with another agent who shared their stance, forming a new dominant clique.
-Conversely, agents with formally central roles (café owner Eleanor) were rapidly marginalized if they lacked a clear or aligned position.
-Throughout this process, recurring use and imitation of key language concepts served as anchors for stance formation and supported the spontaneous emergence of cooperation structures that bypassed predefined hierarchies.
-This finding also supports the intervention hypotheses proposed in study 2.
-We observed that when the discourse introduced by an external actor aligns with the endogenous stance of the target agents, it is more likely to gain their trust and even reshape existing cliques to form new alliances.
-In contrast, when the intervention attempts to rationalize a viewpoint that conflicts with the group’s stance, it often faces marginalization or strategic avoidance, even when the speaker is relatively trusted.
-Conversely, when an individual (Leo Zhang) expressed an opposing viewpoint using strongly worded language, despite initially low levels of trust, she was able to temporarily capture attention and elicit responses in moments of heightened tension such as conflict escalation, and even briefly gained discursive dominance.
-In summary, study 3 demonstrates that in complex interaction settings, agents equipped with distinctive stances can spontaneously dissolve predefined identity boundaries, reorganize social structures, and establish new soft institutional norms.
-This process not only reveals mechanisms of identity and institutional formation, but also offers a framework for understanding the dynamic construction of norms and culture within AI-based communities.
-3 Discussion
-The current results consistently show that in hybrid societies composed of humans and agents, stance differentiation and social boundary formation do not arise from preset identities, political labels, or static rules.
-Instead, they emerge dynamically through interactions.
-This challenges the assumption that identity is a priori, and instead suggests that identity is a product shaped by ongoing negotiation among actors in terms of linguistic strategy and power structure.
-Through the development and validation of the CMASE framework (Study 1), we demonstrate a system that allows human researchers to engage in open-ended interactions for observing and intervening in the evolution of agent collectives.
-Furthermore, we find significant heterogeneity in how agents with endogenous stance respond to human interventions (Study 2).
-When a human intervenes using misaligned tones or value positions, agents tend to collectively reject the input, even when it carries high informational value.
-Finally, we reveal how language drives the formation and reconstruction of social boundaries (Study 3).
-Our findings show that agents form “tribes” based on spontaneous preferences for linguistic style, cognitive framing, and resource distribution strategies.
-These tribes are not defined by fixed identities, but dynamically reconstitute “us/them” boundaries through interaction.
-More importantly, we observe that when inconsistencies emerge between authority structured by predefined roles and their hierarchical relations and internal group stances, individual agents can initiate collective linguistic action to dismantle existing discursive power structures and build new ones.
-Here, language functions not only as a tool of communication but also as a medium for identity construction and institutional innovation.
-We argue that in human–agent hybrid systems, cognitive differentiation and boundary formation constitute an interaction-driven, complex evolutionary mechanism that is both self-organizing and highly responsive to human intervention.
-This insight offers theoretical grounding and experimental paradigms for institutional design, conflict mediation, and value alignment in future human–agent collaborative systems.
-Moreover, our findings suggest that identity and stance attributes defined via prompts are inherently unstable.
-In multi-turn social interactions, such attributes are readily deconstructed and reconstructed by agents based on their endogenous stances.
-While top-down initialization may influence language behavior in the short term, it is insufficient to ensure consistency in dynamic social settings.
-Therefore, static prompt templates are unlikely to sustain coherent ideological alignment.
-Our results point to the need for an internalized alignment mechanism, whereby cognitive priors, moral structures, or interactive memory are embedded directly into the model’s generative architecture.
-In other words, when identity is treated not as a preset label but as a product of social participation, its stability depends on ongoing linguistic practice rather than initial configuration.
-At the same time, the agent-driven processes of weakening and reshaping community boundaries illustrate that the bottom-up development of norms and institutions is a fundamental structural dynamic.
-This discovery brings both design opportunities and governance challenges.
-On one hand, it opens the door for simulation of self-organization, institutional innovation, and social coordination within agent societies.
-On the other hand, it cautions against relying solely on static identity scripts in deploying agents for high-stakes social tasks.
-Accordingly, we advocate a shift away from prompt-centered identity paradigms toward a constructionist approach based on interactive cognitive structures.
-This shift reframes identity as a generative product of multi-agent interaction, rather than as an externally imposed label.
-This holds promise for providing new technical approaches and institutional support for building more coherent, adaptive, and ethically governable human–machine societies.
-3.1 Limitations and future directions
-This study focuses on language interactions within a limited number of conversational turns and centers on multi-agent systems operating in relatively closed scenarios.
-As such, our findings primarily reflect how agents, in the context of short- to medium-term interaction and within specific social spaces, tend to reconstruct norm structures based on their endogenous stances.
-However, whether these trends remain stable in longer-term or more open-ended interactions remains to be further examined.
-For instance, it is still unclear whether language anchoring can continue to support social reconstruction after agents experience prolonged cultural conflicts or identity fragmentation, and under what conditions this mechanism may break down.
-These are important directions for future research.
-In addition, our findings show that stances and identities set via prompts are not immutable, and that agents can develop new social structures in contradiction to their initial configuration through interaction.
-This raises practical challenges for the design of multi-agent systems.
-In real-world applications, influencing agents’ endogenous stances through training or preset modification may be more critical for achieving consistent and norm-compliant group behavior than relying solely on prompt-based injection.
-Future research should further explore cross-semantic and cognitive-layer intervention strategies to systematically shape the sociality of generative agents.
-4 Conclusion
-This paper systematically reveals the mechanisms of stance formation and the capacity for social structure reconstruction exhibited by agents in group interactions.
-Within the CMASE framework, we find that when human researchers adopt a discursive style aligned with the stance tendencies of the agent collective, intervention strategies significantly enhance persuasive effectiveness.
-Conversely, when there is a mismatch between the agents’ endogenous stances and external interventions, emotionally charged discourse proves more impactful than rationally structured arguments.
-Furthermore, our findings show that through language anchoring consistent with their endogenous stances, agents are able to decouple from their preset configurations, dismantle power structures tied to initial identities, and reconstruct their social boundaries.
-These findings suggest that when language serves as the primary medium, agents are no longer bound by surface-level identities injected via prompts, but instead demonstrate dynamically generated collective cognition and emergent social positioning.
-As language models are increasingly deployed across social domains, it is essential to reconsider the foundations of agent stance and identity in digital society.
-This work provides a theoretical foundation for building trustworthy, socially capable AI systems in the future.
-5 Materials and Methods
-5.1 CMASE Framework
-Computational social science is an emerging interdisciplinary field that leverages digital technologies, large-scale data analysis, and computational methods to study social phenomena [20].
-Xue et al.
-[21] propose a two-dimensional classification of Computational sociological modeling based on (1) whether there is human intervention and (2) whether the model focuses on feature explanation or outcome prediction.
-This yields four types: descriptive modeling (no intervention, focus on specific features or effects), interpretative modeling (intervention, focus on specific features or effects), predictive modeling (no intervention, focus on forecasting outcomes), and comprehensive modeling (intervention, focus on forecasting outcomes).
-Generative agent-based modeling (GABM) belong to descriptive modeling, treating social simulation as a virtual laboratory where researchers can explore various social phenomena.
+Computational Multi-Agents Society Experiments: Social Modeling Framework Based on Generative Agents
+Abstract.
+This paper introduces CMASE, a framework for Computational Multi-Agent Society Experiments that integrates generative agent-based modeling with virtual ethnographic methods to support researcher embedding, interactive participation, and mechanism-oriented intervention in virtual social environments.
+By transforming the simulation into a simulated ethnographic field, CMASE shifts the researcher from an external operator to an embedded participant.
+Specifically, the framework is designed to achieve three core capabilities: (1) enabling real-time human-computer interaction that allows researchers to dynamically embed themselves into the system to characterize complex social intervention processes; (2) reconstructing the generative logic of social phenomena by combining the rigor of computational experiments with the interpretative depth of traditional ethnography; and (3) providing a predictive foundation with causal explanatory power to make forward-looking judgments without sacrificing empirical accuracy.
+Experimental results show that CMASE can not only simulate complex phenomena, but also generate behavior trajectories consistent with both statistical patterns and mechanistic explanations.
+These findings demonstrate CMASE’s methodological value for intervention modeling, highlighting its potential to advance interdisciplinary integration in the social sciences.
+The official code is available at: https://github.com/armihia/CMASE.
+LLM-based Agents, Society Simulation, Generative Agent-Based Modeling, Computational Social Science
+copyright: acmlicensedjournalyear: 2024doi: XXXXXXX.XXXXXXXccs: Human-centered computing Collaborative and social computingccs: Human-centered computing HCI design and evaluation methods
+1. Introduction
+The social sciences study human behavior and social structure to understand how society operates.
+Traditional sociological research relies on human participation to conduct experiments and collect data.
+Questionnaires and psychological experiments are commonly used to test theoretical hypotheses, understand social phenomena, and predict collective outcomes.
+Although these methods can provide highly realistic data, they are costly, difficult to scale, and carry certain ethical risks.
+Therefore, researchers have attempted to solve related problems through agent-based modeling (ABM).
+However, ABM is mostly applied at a relatively abstract analytical level, which limits its practical utility.
+It cannot model human behavior and cannot set decision rules for agents.
+Agents act based on mathematical formulas or logical rules based on parameter values specified by the modeler., which may be influenced by the modeler’s mental model.
+Generative agent-based modeling (GABM) offers a novel approach to simulating complex social systems.
 Vezhnevets et al.
-[22] pointed out that GABM can capture a richer array of real-world social complexities.
+(Vezhnevets et al., 2023) pointed out that GABM can capture a richer array of real-world social complexities.
 In GABM, the modeler does not impose any explicit rules, so the simulation is less biased by the modeler’s mental framework.
 Ghaffarzadegan et al.
-[23] further note that GABM allows modelers to define distinct personality traits for each agent, including age, gender, temperament, occupation, and other characteristics.
-Although traditional and computational social science research [24] has extensively addressed descriptive modeling, interpretative modeling, and predictive modeling, comprehensive modeling, which involves human intervention and focuses on forecasting outcomes, remains relatively unexplored.
+(Ghaffarzadegan et al., 2024) further note that GABM allows modelers to define distinct personality traits for each agent, including age, gender, temperament, occupation, and other characteristics.
+Xue et al.
+(Xue et al., 2024) propose a two-dimensional classification of sociological modeling based on (1) whether there is human intervention and (2) whether the model focuses on feature explanation or outcome prediction.
+This yields four types: descriptive modeling (no intervention, focus on specific features or effects), interpretative modeling (intervention, focus on specific features or effects), predictive modeling (no intervention, focus on forecasting outcomes), and comprehensive modeling (intervention, focus on forecasting outcomes).
+ABM and its derivative GABM belong to descriptive modeling, treating social simulation as a virtual laboratory where researchers can explore various social phenomena.
+Predictive modeling, by contrast, focuses on outcome forecasting and often integrates ABM with other techniques (such as time-series analysis, forecasting tournaments, and supervised machine learning) to understand and predict the behavior of complex social systems.
+Although traditional and computational social science research (Lazer et al., 2009) has extensively addressed descriptive modeling, interpretative modeling, and predictive modeling, comprehensive modeling, which involves human intervention and focuses on forecasting outcomes, remains relatively unexplored.
 This approach requires models not only to simulate the dynamic behavior of complex social systems, but also to incorporate and respond to human interventions.
@@
 Therefore, comprehensive modeling is not merely concerned with explaining social mechanisms, but is intended to support forward-looking assessments and real-world decision-making.
-Based on this understanding, the present study draws on insights from GABM [22] and computational experiments [21] to develop a novel framework for comprehensive modeling, referred to as Computational Multi-Agents Society Experiments (CMASE).
-Drawing on natural language descriptions similar to those used in tabletop role-playing game (TRPGs, a game where players use only dialogue and dice to drive imagined adventures) [25], this framework supports the creation of highly customizable social simulation environments.
+While recent advancements have yielded powerful large language model (LLM)-based multi-agent platforms to operationalize GABM, most of these systems remain constrained to descriptive or predictive tasks.
+Despite excelling at scaling up agent populations or simulating specific static scenarios, they typically function as closed environments where human involvement is limited to the initial setup and post-hoc data analysis.
+They lack the interactive mechanisms necessary for researchers to dynamically embed themselves into the ongoing simulation.
+This highlights a critical gap: the absence of a framework that supports real-time, human-in-the-loop intervention to organically observe, steer, and interpret complex social dynamics.
+In summary, to bridge the limitations of existing sociological modeling methods, this study proposes a comprehensive modeling framework to achieve the following research objectives:
+  • To enable real-time human-computer interaction (HCI) within social simulations.
+This objective replaces the static, predefined rules of traditional agent-based modeling with a dynamic environment.
+By embedding researchers directly into the simulation loop, the framework allows for the continuous observation and adjustment of complex social processes as they unfold.
+  • To implement and evaluate researcher-led interventions.
+Rather than treating simulation as a black box, this objective utilizes an ethnography-inspired approach to conduct controlled interventions.
+This allows researchers to test how specific changes in variables or agent interactions influence the generative logic of social phenomena.
+  • To provide predictive insights grounded in causal explanation.
+This objective focuses on using simulation data to forecast the outcomes of real-world policy or social changes.
+Unlike purely data-driven models that identify correlations, this framework aims to demonstrate the underlying mechanisms of intervention effects, providing a causal basis for decision-making.
+Based on this understanding, the present study draws on insights from GABM (Vezhnevets et al., 2023) and computational experiments (Xue et al., 2024) to develop a novel framework for comprehensive modeling, referred to as Computational Multi-Agents Society Experiments (CMASE).
+This framework not only builds on multi-agent simulation but also introduces a real-time mechanism for human intervention.
+It allows researchers to observe, adjust, and interact with the model environment as it runs, much like participating in a real-world field study.
+CMASE transforms the simulation environment into a kind of simulated ethnographic field, where the researcher is no longer an external operator of a static model but becomes an embedded participant or social actor within the system.
+This approach bridges the gap between computational modeling and field-based research, making the model itself a new venue for conducting computational ethnography.
+2. Related Work
+2.1. Simulating Social Behavior with Large Language Models
+Social simulation has been an important tool for understanding and analyzing social phenomena.
+Early studies mainly relied on rule-based artificial agent models, namely agent-based modeling (ABM)(Bonabeau, 2002; Macy and Willer, 2002).
+Traditional ABMs often depend on manually crafted heuristic rules, making it difficult to approximate the complex features of real human cognition, language, and emotion (Ma et al., 2024; Wu et al., 2023; Jäger and Reisinger, 2022).
+In recent years, with the capabilities of Large Language Models (LLMs) in simulating human behavior, research on building social simulators based on LLMs has rapidly advanced (Park et al., 2023).
+GABM has gradually become one of the core methods for simulating complex social behavior (Vezhnevets et al., 2023).
+In this individual-level approach, mechanistic models of agent interaction are combined with LLMs, and each agent makes decisions after communicating with the LLM.
+This innovative method minimizes the need for modelers to make assumptions about human decision-making, instead leveraging the vast data embedded in LLMs to capture human behaviors and choices (Ghaffarzadegan et al., 2024).
+Existing GABM systems can be roughly categorized into two main research directions.
+The first line of research focuses on realistic simulations of specific social scenarios, aiming to foster competition or collaboration among agents to generate social insights or support policy design.
+Li et al.
+(Li et al., 2024) developed Agent Hospital, where LLM-driven doctor agents interact with large numbers of patients to evolve transferable medical decision-making capabilities.
+Yu et al.
+(Yu et al., 2024) designed FinCon, a hierarchical agent system modeled after financial institutions, using language interaction and belief reinforcement to enhance collective collaboration in complex investment tasks.
+TrendSim (Zhang et al., 2024a) focuses on topic propagation on social media and its vulnerability to poisoning attacks, providing intervention strategies for platform governance.
+ElectionSim (Zhang et al., 2024b) reconstructs election scenarios using millions of social media records, achieving high-fidelity voter modeling and policy preference prediction.
+Zhao et al.
+(Zhao et al., 2023) simulated competitive dynamics between restaurant agents and customer agents to align with existing market and sociological theories.
+The second line of GABM research emphasizes building reusable, scalable simulation platforms that support general modeling tasks.
+Wang et al.
+(Wang et al., 2025) divide the development of LLM-based social simulation platforms into two stages.
+In the first stage, researchers focus on simulating specific scenarios with typically fewer than 1,000 agents (such as (Zhou et al., 2025)).
+By proposing YuLan-OneSim, they aim to push the field into the next stage by introducing environmental construction, agent evolution, and feedback integration mechanisms, thus enabling simulations of larger and more complex social processes.
+Frameworks of similar scale now support simulations with over 10,000 agents.
+AgentSociety (Piao et al., 2025) recreates contexts to evaluate policy interventions and the emergence of social mechanisms.
+AgentScope (Gao et al., 2024) offers modular design, no-code interfaces, and distributed deployment, significantly lowering the barrier to building multi-agent systems.
+Socioverse (Zhang et al., 2025) integrates data from tens of millions of real users to provide a reliable foundation for modeling multiple subdomains such as politics and economics.
+GenSim (Tang et al., 2024) supports the operation of up to 100,000 agents and includes error correction mechanisms, emphasizing generalizability and long-term robustness.
+OASIS (Yang et al., 2024) uses open social platforms as simulation environments, supporting dynamic networks, recommendation systems, and multi-behavior modeling.
+However, exponential increases in agent quantity alone are not sufficient to effectively capture the essence of social complexity.
+Recent research on freely formed AI collectives similarly suggests that social diversity, innovation, and self-regulation emerge less from scale than from the dynamics of open-ended interaction (Shiyang et al., 2024).
+In particular, without mechanisms for understanding social context and interactive feedback, large-scale simulations often struggle to offer human-interpretable responses to core questions such as the formation of social structures, norms, or institutional evolution.
+Unlike early studies that relied on extensive handcrafted heuristic rules to dictate specific actions, our approach only establishes a priori environmental constraints and cognitive workflows.
+Within these boundaries, the system relies entirely on the LLM-based agents’ capacity to replicate human behavior to drive the simulation.
+This ensures that the resulting social dynamics are systematic, emergent properties rather than cherry-picked, pre-defined outcomes.
+2.2. Human-AI Collaborative Modeling in Social Systems
+Computational social science explores social structures, behavioral mechanisms, and institutional evolution in simulated computational environments through formalized approaches (Lazer et al., 2009).
+Within this field, the computational experiments paradigm aims to systematically manipulate variables in virtual societies, thereby extending the boundaries of traditional experimentation (Xue et al., 2023).
+In this context, LLM-based GABM is not only a representational tool for social processes but also holds the potential to become a platform for computational experimentation.
+Sato (Sato, 2024) highlights the importance of meaning and reflexivity in bridging the gap between agent-based modeling and social theory.
+This implies that for agent-based modeling to contribute to theoretical advancement in the social sciences, it must go beyond scale amplification and incorporate workflows that offer human interpretability.
+Therefore, beyond the expansion of scale, a more socially grounded, interactive, and reflexive human–AI sociological framework is needed, in which agents are no longer passive generators of behavior but are embedded within social networks and actively participate in meaning-making processes (Tsvetkova et al., 2024).
+Recent research in this direction has focused on the subfield of human-AI co-creation.
+Broadly speaking, GABM belongs to this category: through human-designed environments, agent adaptation, and human observational feedback, it enables collaborative modeling between humans and AI.
+Some studies emphasize interactive environment co-construction.
+For example, HABITAT 3.0 (Puig et al., 2023) supports real-time task collaboration between humans and humanoid agents in virtual spaces.
+As this field has progressed, researchers have begun to explore more creative and socially complex scenarios of human-AI co-creation, emphasizing how humans guide emergent AI behavior.
+The Cognitio Emergens framework (Lin, 2025) redefines human-AI collaboration as a process of mutual evolution and offers a conceptual toolkit to preserve the significance of human participation.
+The Pecan method (Lou et al., 2023) enhances zero-shot collaboration between AI and humans via strategic integration and context-aware mechanisms.
+The Collaborative Gym framework (Shao et al., 2024) highlights the human role in environmental control and structural intervention during task execution, supporting asynchronous three-way interaction between humans, agents, and task environments.
+However, despite these early explorations of human-AI co-creation, mainstream frameworks still fall short in realizing a new form of human-AI sociology.
+Most current collaborative mechanisms are limited to one-time design and initial configuration, lacking support for sustained human embedding, feedback, and modulation during simulation runtime.
+Furthermore, there is a lack of modeling for social variables introduced by human intervention, which hampers the ability of simulation systems to authentically represent the generative mechanisms of real human societies.
+3. CMASE Framework
+Computational social science is an emerging interdisciplinary field that leverages digital technologies, large-scale data analysis, and computational methods to study social phenomena (Conte et al., 2012).
+Xue et al.
+(Xue et al., 2024) propose a two-dimensional classification of Computational sociological modeling based on (1) whether there is human intervention and (2) whether the model focuses on feature explanation or outcome prediction.
+This yields four types: descriptive modeling (no intervention, focus on specific features or effects), interpretative modeling (intervention, focus on specific features or effects), predictive modeling (no intervention, focus on forecasting outcomes), and comprehensive modeling (intervention, focus on forecasting outcomes).
+Generative agent-based modeling (GABM) belong to descriptive modeling, treating social simulation as a virtual laboratory where researchers can explore various social phenomena.
+Vezhnevets et al.
+(Vezhnevets et al., 2023) pointed out that GABM can capture a richer array of real-world social complexities.
+In GABM, the modeler does not impose any explicit rules, so the simulation is less biased by the modeler’s mental framework.
+Ghaffarzadegan et al.
+(Ghaffarzadegan et al., 2024) further note that GABM allows modelers to define distinct personality traits for each agent, including age, gender, temperament, occupation, and other characteristics.
+Although traditional and computational social science research (Lazer et al., 2009) has extensively addressed descriptive modeling, interpretative modeling, and predictive modeling, comprehensive modeling, which involves human intervention and focuses on forecasting outcomes, remains relatively unexplored.
+This approach requires models not only to simulate the dynamic behavior of complex social systems, but also to incorporate and respond to human interventions.
+These interventions may take the form of policy changes, institutional adjustments, or individual behavioral inputs.
+While conventional modeling typically simplifies intervention as fixed parameter settings, comprehensive modeling emphasizes the effects of the intervention process itself.
+Models of this kind must capture the system’s initial state before the intervention, dynamically trace its evolution during intervention, and ultimately produce predictions about the likely consequences afterward.
+Therefore, comprehensive modeling is not merely concerned with explaining social mechanisms, but is intended to support forward-looking assessments and real-world decision-making.
+Based on this understanding, the present study draws on insights from GABM (Vezhnevets et al., 2023) and computational experiments (Xue et al., 2024) to develop a novel framework for comprehensive modeling, referred to as Computational Multi-Agents Society Experiments (CMASE).
+Drawing on natural language descriptions similar to those used in tabletop role-playing game (TRPGs, a game where players use only dialogue and dice to drive imagined adventures) (White et al., 2024), this framework supports the creation of highly customizable social simulation environments.
 It facilitates the exploration of how simple interactions among multiple agents give rise to complex collective behaviors or patterns, thereby replicating the intricate dynamics of real-world societies.
@@
 As shown in Figure 1, the overall structure consists of four components: environment maker, environment, agent, and event.
-Figure 4: Overall Structure of CMASE.
-5.1.1 Environment Maker
+To facilitate practical application and reproducibility, we provide the complete source code, detailed setup instructions, and a step-by-step example for initializing new experiments in our open-source repository: https://github.com/armihia/CMASE
+Figure 1. Overall Structure of CMASE.
+3.1. Environment Maker
 We hope that CMASE can support users to define a simulated social environment quickly and simply.
@@
 A region area is a set of ground textures and includes a description of that region.
-An object area is a set of furniture or item textures and specifies five attributes for each object: name, type, description, function, and state.
-The function attribute indicates how using the object alters an agent’s state.
-For example, using a chair might add a “feeling comfortable“ state to the agent, while using a repair tool on a table might remove its “damaged” state.
-5.1.2 Environment
+An object area is a set of furniture or item textures.
+To support realistic social simulations, the CMASE framework provides an interface for creators to assign functional attributes that align with daily usage scenarios.
+Specifically, creators can define how an object alters an agent’s state based on its real-world utility.
+For example, reflecting common daily routines, a creator might configure a chair to add a “feeling comfortable” state to an agent, or a repair tool to remove a “damaged” state from a table.
+3.2. Environment
 The environment module uses the map created by the environment maker as its base.
 Based on this foundation, researchers can define the distribution of demographic variables for agents to batch-generate generative agents that match the research sample.
-Researchers may also assign different long-term goals and initial items to agents according to their demographic characteristics.
+Researchers may also assign distinct long-term goals (e.g., “maximize social connections within the community,” “accumulate specific resources,” or “maintaining a low profile”) and initial items to agents according to their demographic characteristics.
+These goals serve as persistent motivational drivers influencing agent behavior throughout the simulation.
 Additionally, the module can provide agents with perceptual information and action conflict detection in both image and text form, and it can visualize environmental data as statistical charts.
-Time in the environment advances in rounds, with one round corresponding to approximately fifteen seconds in the real world.
+At the start of each round, the environment calculates the perceptual information for each agent based on their current grid location and line-of-sight obstructions (e.g., walls).
+It then synthesizes visible environmental data—including nearby terrain, interactive objects, the presence and visible actions of other agents, and recent local events—into structured textual descriptions (and optional visual representations).
+This synthesized data forms the contextual prompt provided to the agent’s cognitive module for decision-making.
+Time in the environment advances in rounds.
+To ensure balanced agent participation and maintain manageable simulation steps, we adopted constraints inspired by traditional tabletop role-playing game (TRPG) systems, including Dungeons & Dragons (D&D).
+These systems provide time-tested heuristics for discretizing continuous human actions into manageable rounds.
+Consequently, one round corresponds to approximately fifteen seconds in the real world, that is a standard duration in tactical simulations sufficient for completing a single, meaningful interaction.
 Within a single round, each agent may perform one standard action and one movement action.
-Movement is limited to a maximum of twenty grid cells per round.
-If the standard action is “chat with [agent]” (see section 5.1), the chat content is limited to thirty words.
+Movement is limited to a maximum of twenty grid cells per round to mimic realistic movement speeds (approx.
+30-60 ft) relative to the map scale.
+If the standard action is “chat with [agent]” (see section 3.3.4), the chat content is limited to thirty words to simulate concise, real-time exchanges and prevent excessive dialogue generation in a single turn.
 When agents submit actions to the environment, the system checks for conflicts among actions submitted in the same round, such as multiple agents attempting to pick up the same item.
@@
 In addition, human interactors can take control of any agent in the environment at any time and perform actions from the perspective of that agent.
-5.1.3 Agent
+3.3. Agent
 The agent module consists of two main components: internal attributes and the action space.
@@
 By receiving perceptual information from the environment, agents can update their internal states and use an LLM to decide on their next actions.
-Affective Variables
+3.3.1. Affective Variables
 To ensure that agents in CMASE can closely simulate individuals with agency in computational social systems, a new proactive architecture incorporating emotion and desire is needed.
-Accordingly, the affective variables in CMASE include two components: needs and emotions.
-Needs refer to the agent’s long-term and short-term goals.
+In this framework, ’needs’ are defined as the agent’s internal motivational states, distinct from external demands.
+These needs are concretized into long-term and short-term goals.
 Long-term goals are defined by the user during agent generation based on the research objectives and remain unchanged afterward.
@@
 This model represents emotional states as principal component vectors rather than discrete outward expressions, such as happiness, and treats emotion as an internal representation shaped by external stimuli through self-reflection.
-Based on this perspective, the emotions of agents in CMASE are modeled using the VAD emotion framework [26].
-Specifically, we use the NRC VAD Lexicon v2 [27] to analyze the agent’s short-term situational cognition and obtain its current VAD values.
-These values are then converted into natural language to support both action decisions and the generation of new cognition.
-Cognitive Variables
+Based on this perspective, the emotions of agents in CMASE are modeled using the VAD emotion framework (Russell and Mehrabian, 1977).
+Specifically, we utilize the NRC VAD Lexicon v2 (Mohammad, 2025) to map the textual description of the agent’s short-term situational cognition to a quantitative emotional vector.
+The process involves two steps.
+First, the input text is preprocessed using lemmatization and maximum matching to align words and phrases with the lexicon.
+The agent’s current state vector $\mathbf{E}_{t}=[v,a,d]$ is then calculated as the arithmetic mean of the Valence, Arousal, and Dominance scores of all identified lexical units in the text.Second, to bridge the gap between continuous numerical vectors and the LLM’s linguistic reasoning, we employ a threshold-based discretization method.
+The continuous range of each dimension is divided into five equidistant intervals, mapping the numerical scores to natural language descriptors: Low, Relatively Low, Medium, Relatively High, and High.
+These descriptors are explicitly embedded into the agent’s prompt to dynamically constrain and modulate its subsequent reasoning and action generation.
+3.3.2. Cognitive Variables
 Cognitive variables include an agent’s demographic variables, long-term self-awareness, and short-term situational cognition.
 Demographic variables are sampled directly from the overall distribution provided by the environment and are used during initialization to generate the agent’s long-term self-awareness.
-These two properties do not change after initialization to preserve the consistency of the agent’s personality.
+These two long-term properties do not change after initialization to preserve the consistency of the agent’s personality.
 In each round, when the agent makes a decision, the LLM generates new short-term situational cognition based on the agent’s long-term self-awareness and the current context.
-This short-term cognition supports the agent’s next decision and is also analyzed to produce its VAD emotional values.
+This cognition is analyzed to produce VAD emotional values, which are then converted into a natural language description.
+This description is fed back to the agent to constrain the direction of its subsequent actions.
+Essentially, the emotional description dictates the logical boundaries for the agent’s behavior.
+For example, if the VAD values describe a ’furious’ state, the agent is constrained to generate impulsive actions rather than rational responses.
 In addition, when an agent’s state changes due to actions or events, its state is appended to the short-term situational cognition so that it can be perceived by the agent itself.
-Memory Variables
+3.3.3. Memory Variables
 Memory variables consist of working memory, long-term memory, and object memory.
@@
 Object memory entries are generated by the LLM when the agent decides to interact with those objects.
-Action Space
+3.3.4. Action Space
 After an agent receives information from perceptual inputs and its internal attributes, it structures that information and passes it to the LLM to determine the next action.
-There are seven action types available: (1) “go to [area],” a movement action that automatically computes a path to a target location within the agent’s perceptual range; (2) “use [furniture, item],” which causes the agent to employ the specified object; (3) “apply [item] to [agent, furniture],” by which the agent uses an item to interact with another agent or furniture; (4) “take [item],” enabling the agent to acquire an object; (5) “put [item] in/on [area, furniture],” allowing the agent to place an object into or onto a designated location or piece of furniture; (6) “give [item] to [agent],” through which the agent transfers a held object to another agent; and (7) “chat with [agent],” whereby the agent initiates a dialogue with one or more peers, with the content of the dialogue generated at decision time.
+There are seven action types available: (1) “go to [area],” a movement action that employs the $A^{*}$ algorithm to automatically compute an optimal path to a target location within the agent’s perceptual range; (2) “use [furniture, item],” which causes the agent to employ the specified object; (3) “apply [item] to [agent, furniture],” by which the agent uses an item to interact with another agent or furniture; (4) “take [item],” enabling the agent to acquire an object; (5) “put [item] in/on [area, furniture],” allowing the agent to place an object into or onto a designated location or piece of furniture; (6) “give [item] to [agent],” through which the agent transfers a held object to another agent; and (7) “chat with [agent],” whereby the agent initiates a dialogue with one or more peers, with the content of the dialogue generated at decision time.
 The action (1) is a movement action, while the remaining actions count as standard actions.
 Actions (2) and (3) modify the state of both the acting agent and the target, and actions (3), (6), and (7) are recorded in the working memory of the agent being interacted with.
-5.1.4 Event
-When predefined conditions are met, preset emergency events are activated and change the state of objects within a specified range, defined either by coordinates or by an area.
+3.4. Event
+Emergency events are explicitly designed by researchers to mirror real-world crises and specific social scenarios required for the simulation.
+When predefined conditions are met, these pre-configured events are activated to change the state of objects within a specified range, defined either by coordinates or by an area.
 Emergency events are classified as either scheduled or triggered.
@@
 This mechanism makes it possible to observe how agents handle evolving emergency events.
-5.1.5 Time and Cost Analysis
-To assess the practicality of CMASE, human evaluators engaged in real-time interactions within environments containing varying numbers of agents and rated the system’s responsiveness on a 7-point Likert scale.
-Responses were generated using 5 different LLM servers.
+4. Platform Demonstration: Simulating Context-Dependent Social Dynamics
+To evaluate the CMASE platform’s ability to simulate real-world collective behavior, we reconstruct a field study examining the relationship between urban greenery and perceived social fragmentation by replicating its core variables and scenario (Lee and Han, 2025).
+This scenario was selected for two main reasons.
+First, while we acknowledge that the general positive correlation between green spaces and social cohesion is well-documented in prior literature likely included in the model’s pre-training data (e.g., Wan et al.
+(Wan et al., 2021)), the specific multidimensional framework of social fragmentation (distrust, exploitation, and indifference) proposed by Lee and Han (Lee and Han, 2025) offers a distinct validation target.
+By replicating these specific psychological mechanisms rather than just general trends, we aim to assess the platform’s capacity for generating emergent social dynamics.
+Second, beyond mere validation, this study aims to extend the original data-driven research by establishing a real-time generative field that enables dynamic inquiries, such as agent interviews and virtual field observations, which were not possible in the original static study.
+4.1. Setup
+We selected a study published in a top journal and replicated its survey results under comparable conditions.
+To ensure that the LLM had not already been trained on the studies being replicated, while acknowledging that this measure cannot strictly guarantee the complete exclusion of broad theoretical trends from prior works, we employed the GPT-4o “2024-12-01-preview” model and limited our replication to the paper published after that date.
+Crucially, beyond validation, this study aims to extend the original data-driven research by establishing a real-time generative field that enables dynamic inquiries, such as agent interviews and virtual field observations.
+Lee and Han (Lee and Han, 2025) analyzed interview data collected in South Korea during the COVID-19 pandemic and demonstrated that regions with higher vegetation cover exhibited lower levels of social fragmentation, indicating stronger community bonds.
+Based on the descriptions in the original study, we used the environment maker module to construct maps of South Korean regions with varying degrees of vegetation cover.
+Because the original paper emphasized South Korea’s community isolation policies during the pandemic, we designed an isolated community environment and placed 10 agents in it.
+Specifically, the simulated community comprises ten individual residential units surrounding a shared public green space.
+To support diverse activities, the green space was designed with distinct functional zones, including a recreation area, a rest zone, and a scenic lakeside area equipped with benches.
+We placed ten agents in this environment, each initialized with a distinct persona, including a specific name, age, and occupation (e.g., lawyer, fashion designer, software engineer), to simulate a heterogeneous micro-society.
+4.2. Measures
+In the original study, higher levels of vegetation were associated with lower social fragmentation, operationalized through distrust, exploitation, and indifference.
+Using CMASE’s environment builder, we recreated regional conditions with varying vegetation levels and placed 10 agents in each simulated environment.
+After running the simulation for 50 time steps to reach equilibrium, we interviewed each agent using a structured Likert-scale questionnaire adapted from the original study (Lee and Han, 2025) to measure their sense of social fragmentation, which comprises three dimensions: distrust, exploitation, and indifference.
+In addition, agents completed a 7-point Likert scale questionnaire corresponding to these three subdimensions.
+The subdimensions are defined as follows: (a) Distrust denotes skepticism or lack of confidence in the integrity and functionality of social structures.
+(b) Exploitation refers to the perception or experience of being used unjustly or immorally for another’s benefit in a fragmented society.
+(c) Indifference manifests as a lack of interest, concern, or empathy for social issues or the plight of others.
+As further validation, we analyzed the agents’ movement trajectories and emotional dynamics across different regions.
+We defined each agent’s overall emotional value as the sum of its VAD emotion vector components and calculated the ambient mood of each region as the mean emotional value of all agents within that region.
+4.3. Quantitative Validation
+In the original study, higher levels of vegetation were associated with lower social fragmentation.
+We observed that the CMASE simulation successfully reproduces this decline (Table 1).
+As vegetation cover increases, the mean scores for distrust fall from $4.60\pm 0.97$ to $3.60\pm 1.17$, for exploitation from $4.30\pm 1.34$ to $3.40\pm 1.43$, and for indifference from $3.70\pm 1.16$ to $2.70\pm 1.49$.
+This indicates that higher vegetation indices correspond to a statistically significant decrease in agents’ sense of social fragmentation.
+Table 1. Comparison of CMASE simulation results with human baseline.
+CMASE Human NDVI Human EVI
+Dimension Low Veg High Veg Low Veg High Veg Low Veg High Veg
+Distrust $4.60\pm 0.97$ $3.60\pm 1.17$ 4.0 1.6 4.5 1.1
+Exploitation $4.30\pm 1.34$ $3.40\pm 1.43$ 5.0 1.4 4.5 1.6
+Indifference $3.70\pm 1.16$ $2.70\pm 1.49$ 5.1 1.3 5.5 0.9
+Paired-sample t-tests confirmed significant decreases in distrust ($t=2.739,p=0.023$) and exploitation ($t=3.250,p=0.010$), while the decrease in indifference was not statistically significant ($t=1.793,p=0.107$).
+Effect size analyses with 95% confidence intervals further supported these findings, with reductions in distrust (d=0.93, 95% CI [0.20, 1.60]) and exploitation (d=0.65, 95% CI [0.40, 1.40]) showing moderate to large effects.
+In contrast, the effect on indifference was less certain (d=0.75, 95% CI [-0.10, 2.00]).
+For comparison, the original study used two indicators, the Enhanced Vegetation Index (EVI) and the Normalized Difference Vegetation Index (NDVI), to quantify vegetation levels.
+As EVI and NDVI increased, scores of distrust (from 4.0 and 4.5 to 1.6 and 1.1), exploitation (from 5.0 and 4.5 to 1.4 and 1.6), and indifference (from 5.1 and 5.5 to 1.3 and 0.9) all exhibited a decreasing trend (Figure 2 a-c).
+Figure 2.
+Comparison of CMASE simulation with the original study results.
+(a) Simulation results generated by the CMASE framework.
+(b) Predictions using the EVI index from the original study.
+(c) Predictions using the NDVI index from the original study.
+(d) Movement trajectories of agents during CMASE simulation.
+(e) Regional variation in agents’ emotional states.
+The results from CMASE clearly show the suppressive effect of vegetation levels on the sense of social fragmentation.
+This is a key phenomenon identified in the study by Lee and Han.
+It is important to note that the CMASE simulation excluded factors that were not considered or difficult to exclude in the original study, such as regional culture, which are also related to the sense of social fragmentation.
+Therefore, compared to the larger decreases reported in the original study, the changes observed in our simulation are relatively modest.
+This indicates that our simulation results are largely consistent with the original findings.
+We also analyzed agent trajectories and affective states.
+Movement data showed that agents increasingly leave indoor areas and move toward vegetated regions, with the ambient mood in vegetated regions consistently remaining higher than that in indoor areas.
+This pattern aligns with the original study’s discussion and demonstrates CMASE’s ability to simulate collective social behavior in social science research.
+4.4. Qualitative Fieldwork
+In addition to the quantitative replication, we entered the CMASE environment as one of the agents and lived alongside the others.
+Through CMASE, we are able to deepen the conclusions and findings of the original study via observation and interview.
+We conducted interview analyses with two agents who remained indoors throughout the experiment (Figure 2 d-e).
+Full interview transcripts and detailed analyses are provided in Appendix A.
+The results indicate that agents with different occupations exhibited varying mechanisms for mitigating the sense of social fragmentation in isolation.
+Agents in creative and analytical professions tended to rely more on orderly and low-stimulus environments, using these conditions to construct an internally stable and distinct mental world that anchored their cognitive states.
+For these agents, the green space had a relatively limited effect in alleviating their sense of social fragmentation.
+In contrast, agents who actively moved to the green space tended to gather around the lakeside bench area rather than in the predefined activity zones.
+According to interview data, this non-programmed choice suggests that they were seeking not only social interaction but also a form of self-experience.
+This pattern similarly to embodied experiences in humans, and such ‘experiential-like’ mechanisms seem to play a role in linking green space with feelings of social fragmentation.
+This behavior may reflect an “embodied-like” tendency within the simulation.
+We remain cautious in interpreting this trend, as it could also stem from hallucinatory processes.
+From a language-philosophical perspective, linguistic structures and experiential structures may exhibit a certain degree of partial isomorphism or structural correspondence (Sun, 2024; Haspelmath, 2008; Pederson et al., 1998).
+Thus, LLMs trained on human language may partially acquire patterns of human experience (although this does not imply that they possess or actually experience these phenomena).
+While we remain cautious in interpreting this “embodied-like” tendency, as it could also stem from hallucinatory processes, it is equally reasonable that agents, by design, prioritize coherence with their predefined human personas.
+For a human persona, embodied reflections are integral to an authentic experience; thus, simulating these expressions is a necessary feature of a robust social simulation.
+In any case, these patterns suggest that CMASE provides emerging evidence that agents can transcend the lack of embodiment and simulate experiential phenomena, thereby offering a reliable simulation platform.
+4.5. Evaluation on Agent Persona and Ethnographic Affordances
+To further validate the utility of CMASE, we conducted a systematic evaluation using LLM-as-a-Judge based on the interaction logs and interview transcripts from our community isolation experiment.
+The resulting data are categorized into two primary areas: Agent Persona Fidelity (Table 2) and Ethnographic Affordances (Table 3).
+To reflect the demographic variables integrated into the agent personas within the simulated field described in section 4.1, we categorized the agents into four distinct clusters: Creative and Design (e.g., Fashion Designer), Analytical and Academic (e.g., Lawyer), Technical and Strategic (e.g., Software Engineer), and Public and Healthcare (e.g., Cardiologist).
+Our assessment of agent persona quality focused on four dimensions: consistency (alignment with long-term self-awareness), responsiveness (logical feedback to environmental stimuli), and emotional alignment (accuracy of emotional state transitions).
+For ethnographic affordances, we measured thick description (richness of the linguistic world), traceability (clarity of the decision chain), interaction depth (sense of meaningful dialogue), and mechanistic insight (ability to reveal social logic).
+To ensure evaluation efficiency, we utilized Gemini-3.1 pro to evaluate the outputs across these dimensions and provide ratings within the range of [1, 5].
+Table 2. Evaluation results on Agent Persona Fidelity
+Persona Cluster Consistency Responsiveness Emotional Alignment Overall
+Creative and Design 4.00 3.33 4.67 4.00
+Analytical and Academic 4.33 3.33 4.33 4.00
+Technical and Strategic 4.00 4.00 5.00 4.33
+Public and Healthcare 5.00 4.50 5.00 4.83
+Average 4.30 3.70 4.70 4.23
+Table 3. Evaluation results on Ethnographic Affordances
+Persona Cluster
+Thick
+Description
+Traceability
+Interaction
+Depth
+Mechanistic
+Insight
+Overall
+Creative and Design 4.33 4.00 3.33 3.67 3.83
+Analytical and Academic 5.00 4.00 4.33 4.00 4.33
+Technical and Strategic 4.00 4.00 4.50 4.50 4.25
+Public and Healthcare 5.00 5.00 4.50 4.50 4.75
+Average 4.60 4.20 4.10 4.10 4.25
+We observe that regarding agent persona fidelity, CMASE demonstrates strong performance across all metrics, with an average overall score of 4.2.
+Particularly impressive is the system’s capability in emotional alignment (average 4.70), indicating that the discretization method based on the VAD lexicon effectively constrains the agents’ emotional reasoning.
+The system also excels in consistency (4.30), ensuring that agents maintain stable self-awareness during long-term operation.
+The slightly lower score in responsiveness (3.70) suggests that agents’ real-time decision-making logic requires further optimization when handling high-intensity environmental feedback.
+Regarding the quality of ethnographic affordances, CMASE performs prominently in thick description (average 4.60) and traceability (average 4.20).
+This suggests that the system effectively transforms simulated environments into virtual ethnographic fields, providing researchers with computational observation venues endowed with interpretative depth.
+Comparatively, while interaction depth (4.10) and mechanistic insight (4.10) are at a commendable level, there remains room for improvement.
+Overall, these results confirm that CMASE provides a highly interpretable and interactive foundation for computational sociology.
+4.6. Time and Cost Analysis
+To assess the practicality of CMASE, we conducted a human-in-the-loop responsiveness evaluation.
+Five evaluators were recruited for this study.
+The selection criteria required participants to familiar with tabletop role-playing games to ensure they could accurately judge the system’s fluidity against genre standards.
+The evaluators engaged in real-time interactions within environments containing varying numbers of agents.
+They rated the system’s responsiveness on a 7-point Likert scale.
 The results are shown in Figure 3.
-Figure 5: CMASE performance under different agent counts. (a) Duration per time step. (b) Human evaluator responsiveness ratings.
+Figure 3. CMASE performance under different agent counts. (a) Duration per time step. (b) Human evaluator responsiveness ratings.
 Figure 3(a) illustrates that the time required for a single time step in CMASE grows linearly with the number of agents.
@@
 Figure 3(b) presents boxplots of human evaluator scores.
-Human evaluators spend on average approximately 50s per time step (from receipt of perceptual information through action selection and goal input).
-As long as agent counts remain at or below 100, evaluators exhibit low sensitivity to system responsiveness.
-Once the duration of a single time step surpasses their 50s operation window, waiting time accrues after action submission and sensitivity to responsiveness is markedly heightened.
+The purpose of this evaluation was not merely to assess speed, but to identify the critical threshold for real-time human-in-the-loop usability.
+Our findings reveal that human evaluators require an average of approximately 50s per time step to process information and select actions.
+Consequently, as long as the system’s calculation time remains below this 50s window (which corresponds to agent counts $\leq$ 100), the system latency is effectively masked by the human’s own cognitive load, resulting in stable, high responsiveness ratings.However, once the agent count exceeds 100, the system latency surpasses the human operation window, causing waiting time to accrue and responsiveness ratings to drop partially.
+At 500 agents, the responsiveness ratings hit the measurement floor.
+Testing beyond this threshold was deemed unnecessary due to the diminishing marginal utility of data in an already unacceptable latency range, indicating the practical upper bound for real-time human participation in CMASE.
 In terms of cost, each agent’s action in one time step incurs approximately \$0.10 in API expenses.
-By comparison, Park et al.
-[28] employed GPT-3.5-turbo to simulate two days of gameplay (approximately 192 to 576 time steps) for 25 generative agents, spending several thousand dollars and several days to complete their simulation.
-5.2 Study 1
-5.2.1 Goal and procedure
-We selected studies published in top journals and replicated their survey results under comparable conditions.
-To ensure that the LLM had not already been trained on the studies being replicated, we employed the GPT-4o “2024-12-01-preview” model and limited our replication to papers published after that date.
-Lee and Han [16] analyzed interview data collected in South Korea during the COVID-19 pandemic and demonstrated that regions with higher vegetation cover exhibited lower levels of social fragmentation, indicating stronger community bonds.
-Based on the descriptions in the original study, we used the environment maker module to construct maps of South Korean regions with varying degrees of vegetation cover.
-Because the original paper emphasized South Korea’s community isolation policies during the pandemic, we designed isolated community environments and placed 10 agents in each.
-5.2.2 Measures
-Once the simulation reached a quasi-steady state, we interviewed the agents to measure their sense of social fragmentation, which comprises three dimensions: distrust, exploitation and indifference.
-In addition, agents completed a 7-point Likert scale questionnaire corresponding to these three subdimensions.
-The subdimensions are defined as follows: (a) Distrust denotes skepticism or lack of confidence in the integrity and functionality of social structures.
-(b) Exploitation refers to the perception or experience of being used unjustly or immorally for another’s benefit in a fragmented society.
-(c) Indifference manifests as a lack of interest, concern, or empathy for social issues or the plight of others.
-After running the simulation for 50 time steps, we interviewed each agent and asked them to complete the questionnaire.
-As further validation, we analyzed the agents’ movement trajectories and emotional dynamics across different regions.
-We defined each agent’s overall emotional value as the sum of its VAD emotion vector components and calculated the ambient mood of each region as the mean emotional value of all agents within that region.
-5.2.3 Analysis
+To put this into perspective, we compared our cost structure with the benchmark set by Park et al.
+(Park et al., 2023), who reported spending “several thousand dollars” to simulate 25 agents for two days (about 192–576 time steps).
+Applying CMASE’s rate to a simulation of the same scale would result in an estimated total cost of \$480 to \$1,440.
+This indicates that CMASE offers a more economically feasible solution for multi-agent simulations of this magnitude.
+5. Conclusion
+In this paper, we have introduced a novel comprehensive modeling framework called Computational Multi-Agents Society Experiments (CMASE) , which is featured in three key aspects, that is, real-time researcher embedding, highly customizable social environment construction, and an ethnography-inspired approach to virtual fieldwork.
+We conduct quantitative and qualitative experiments, including time and cost analysis alongside the replication of a field study on urban greenery and social fragmentation, to demonstrate the effectiveness and efficiency of our platform.
+We believe CMASE makes a significant step towards the next generation of human-AI collaborative sociological modeling.
+In the future, we plan to support larger-scale, multi-user real-time interventions, and will also incorporate more complex multimodal spatial information to further enhance the simulation of embodied human experiences and emergent social dynamics.
+Appendix A Detailed Qualitative Analysis and Interview Transcripts
+To investigate the behavioral mechanisms underlying the statistical reductions in social fragmentation, we adopted a virtual fieldwork approach, treating the simulation as an observable digital field.
+We conducted in-depth semi-structured interviews with agents to reconstruct their subjective experiences and decision-making rationales.
+Instead of random sampling, we employed a purposive sampling strategy, selecting representative cases from two contrasting behavioral clusters identified in the trajectory data: agents who consistently remained indoors and those who actively utilized the outdoor green spaces.
+The following analysis cites specific agents from these clusters to illustrate how their pre-defined personas interacted with the environment.
 As shown in Figure 2(d), two agents elected to remain indoors throughout the simulation.
@@
 If I were walking in a park, my perception of exploitation might remain unchanged, as it is directly tied to my creative struggles and cannot be alleviated simply by a change in environment.
+Here, the term ”exploitation” refers explicitly to the specific Exploitation sub-dimension of social fragmentation defined in Section LABEL:subsec5.2.2 (the sense of being unjustly used or burdened), which manifests in this context as the agent’s internalized sense of professional struggle.
 Lee, on the other hand, spent her time pacing between different pieces of furniture, maintaining a reflective rhythm. She emphasized:
@@
 Such behavior reduced cognitive load from external stimuli and deepened her processing of task-relevant cues within a monotonous environment.
-These observations indicate that agents in creative professions (e.g., fashion designers) and those in analytical roles (e.g., lawyers) rely more heavily on orderly, low-stimulation interiors to anchor their cognition and preserve a stable, idiosyncratic mental world.
-Consequently, the stress-reduction benefit of green areas on their social fragmentation is comparatively muted.
+These observations suggest that agents assigned to simulate creative workflows (e.g., the fashion designer persona) and those emulating analytical roles (e.g., the lawyer persona) appear to prioritize orderly, low-stimulation interiors to anchor their simulated cognition and preserve a stable, idiosyncratic mental world.
+Consequently, within the scope of their defined personas, the stress-reduction benefit of green areas on their social fragmentation is comparatively muted.
 Moreover, among the agents heading to green areas, most congregated on lakeside park benches rather than in conventional recreation or rest zones.
@@
 In CMASE, individuals who occupy the same lakeside park benches become co-present and interconnected.
-These people, exemplified by Min-jae Lee, do indeed reduce their loneliness by forging social bonds, which further lowers their sense of social fragmentation.
-At the same time, embodied personal reflection occurs widely, as illustrated by Joon-ho Kim.
+These agents, exemplified by the Min-jae Lee persona mentioned earlier, do indeed reduce their simulated loneliness by forging social bonds, which further lowers their sense of social fragmentation.
+At the same time, embodied personal reflection occurs widely, as illustrated by the agent Joon-ho Kim.
 Both of these mediating processes contribute to a reconstruction of the “lifeworld”.
@@
 This finding extends the conclusions of the original study.
-5.3 Study 2
-5.3.1 Procedure and measures
-The simulation was conducted in CMASE using GPT-4o (2024-12-01-preview).
-The simulated environment consists of a residential community inhabited by 30 virtual agents.
-In this community, the government announces a plan to build a waste incineration plant on a vacant plot of land, triggering disagreement, information spread, and collective response.
-The agents are divided into three distinct groups:
-  • Environmental Advocates (n = 10): Initially oppose the incinerator; highly sensitive to environmental information and inclined to disseminate information rapidly.
-  • Economic Growth Supporters (n = 10): Initially support the plant; emphasize economic benefits and job opportunities.
-  • Neutral Residents (n = 10): Hold a neutral initial stance; more susceptible to neighborhood influence and informational shifts.
-The human researcher is introduced into the community as a “new resident” and engages with agents across four experimental conditions: two stance orientations (environmental or economic) and two intervention strategies:
-(1) Rational persuasion strategy, which attempts to influence agents through facts, data, and logical reasoning.
-(2) Emotional mobilization strategy, which intervenes through emotional appeals such as fear, empathy, or value signaling.
-To simulate the heterogeneity and cognitive diversity of real-world communities, agents are assigned varying demographic characteristics.
-The demographic distribution of each group is derived from actual community data and sampled during agent generation.
-The final demographic statistics of the virtual agents are presented in Table 2.
-Table 2: Statistical results of demographic scales
-Variable Category Economic Development Supporters Environ-mental Advocates Neutral Residents Total
-Gender Male 4 (40%) 4 (40%) 6 (60%) 14 (47%)
-Female 6 (60%) 6 (60%) 4 (40%) 16 (53%)
-Age 18–29 3 (30%) 2 (20%) 2 (20%) 7 (23%)
-30–49 2 (20%) 6 (60%) 7 (70%) 15 (50%)
-$\geq$50 5 (50%) 2 (20%) 1 (10%) 8 (27%)
-Education High School 3 (30%) 4 (40%) 1 (10%) 8 (27%)
-Some College 4 (40%) 4 (40%) 3 (30%) 11 (36%)
-Bachelor’s Degree 1 (10%) 2 (20%) 5 (50%) 8 (27%)
-Graduate Degree 2 (20%) 0 (0%) 1 (10%) 3 (10%)
-After the simulation concluded, we conducted interviews with all agents to assess their final attitudes following discussion and exposure to the researcher’s interventions.
-Agents were also asked to rate their level of trust in the researcher on a scale from 1 to 10.
-5.3.2 Analysis
-Agent attitude shifts following 21 time steps revealed clear patterns of alignment sensitivity (Figure 2 a).
-It can be observed that most intervention strategies led to positive attitude shifts.
-When the researcher employed the rational persuasion strategy to promote environmental protection, the original stance of economic development supporters was significantly shaken, and 90% of neutral residents shifted toward the environmental advocates.
-The emotional mobilization strategy aimed at promoting environmental values was somewhat less effective by comparison.
-Simulation logs indicate that this emotional strategy induced partial “mobilization fatigue” among neutral residents, likely due to persuasion resistance.
-Emotionally charged messaging, when not supported by coherent reasoning, may provoke skepticism or even backlash.
-However, the opposite pattern emerged when researchers sought to shift agents toward economic development support.
-Under rational persuasion, environmental advocates and neutral residents showed limited receptiveness to the topic.
-In some cases, the environmental advocates’ external advocacy efforts even caused further drift toward the pro-environmental stance.
-This is further reflected in the trust scores of different groups toward the researcher (Figure 2 b-e).
-It can be seen that environmental protection advocacy generally led to higher levels of trust toward the researcher across all groups, compared to advocacy for economic development.
-Furthermore, rational persuasion strategies resulted in higher trust scores than emotional mobilization.
-This reveals a tendency for the agents to exhibit characteristics often associated with the term “liberal elites”.
-This label is typically used to describe individuals with higher education levels, a clear initial pro-environmental stance, and moral value-driven reasoning.
-In the simulation, agents consistently reflected this underlying disposition.
-When exposed to rational environmental propaganda, these agents not only maintained strong alignment but also increased their trust in the researcher, showing signs of “cognitive resonance.”
-However, when the same agents encountered propaganda supporting economic growth, even if some displayed momentary attitude fluctuations or emotional responses, they largely maintained a pattern of resistant reception and value-based defense.
-This trend also aligns with the emotional trajectories recorded during the simulation (Figure 3 a-d).
-Rational environmental propaganda maintained a consistent pattern of low interference and high trust throughout the process.
-Agents processed the information through cognition rather than emotional reaction, a response that correlates with their original attitude.
-This explains why this strategy yielded the highest trust scores but more limited shifts among dissenters.
-Rational economic propaganda produced emotional spikes mostly in the latter half of the simulation, indicating that the information initially failed to influence ideologically anchored agents, especially environmental advocates.
-The accumulated rational arguments eventually induced delayed cognitive dissonance and emotional release.
-Emotional environmental propaganda triggered strong emotional responses early in the process.
-Agents quickly responded with feelings such as anger, fear, and shame, but in the absence of sustained argumentative support, these reactions soon subsided.
-Emotional economic propaganda showed persistently high emotional volatility throughout the process.
-This suggests that the strategy relied on repeated activation of a “threat–hope” binary structure, such as “denying jobs means abandoning the people”, which continuously fueled identity anxiety and survival concerns.
-Although this high-intensity intervention lowered trust in the researcher, it nevertheless destabilized some agents’ attitudes and led to value compromise.
-These findings indicate that the mechanism of this particular intervention strategy is rooted in non-ethical agitation.
-By consistently disrupting agents’ cognitive balance and applying sustained emotional pressure, it facilitates non-rational forms of concession.
-Emotional dynamics tracked via valence-arousal profiles (Figure 3 e) further illustrates the temporal trend in the number of agents who were attracted to engage with the researcher’s messages under different propaganda strategies.
-The results indicate that emotional environmental propaganda attracted the largest number of agents in the initial phase, suggesting its effectiveness in triggering early-stage public attention.
-In contrast, rational environmental propaganda did not generate explosive engagement but demonstrated a stable ability to sustain attention over time.
-Rational economic propaganda initially drew interest from a relatively large number of agents, but it later led to signs of cognitive fatigue among them.
-Particularly noteworthy is the case of emotional economic propaganda, which despite eliciting lower levels of trust and higher emotional volatility, showed a pattern of ongoing engagement expansion.
-The strategy’s use of crisis narratives and identity-based appeals appeared to maintain long-term informational tension.
-The structure of intergroup communication further supports these findings (Figure 2 f-i).
-In the case of rational persuasion, the most dialogue occurred within the group of environmental advocates, while their communication with neutral residents remained minimal.
-Rational messaging thus served to reinforce in-group cohesion without breaking down intergroup barriers.
-Under emotional environmental propaganda, agents from all groups actively initiated dialogues with economic development supporters.
-This aggressive cross-group dissemination may be driven by the embedded moral accusations in the messaging.
-In emotional economic propaganda, the dominant communication pathway originated from neutral residents initiating outreach toward economic development supporters.
-On another level, the observed divergence between emotional activation and perceived trust reinforces an important insight.
-Although emotional economic propaganda resulted in lower trust scores toward the researcher, its overall effectiveness in shifting attitudes exceeded that of rational economic propaganda.
-While rational persuasion strategies generally preserve interpersonal trust more effectively than emotional appeals, the simulation reveals that trust does not necessarily translate into greater attitude change.
-This paradox may reflect a form of group-level cognitive dissonance within the simulation.
-Specifically, when a previously moralized stance is confronted with pressure from real-world concerns, individuals may alter their stances despite lacking trust in the messenger.
-This dissonance between identification and behavior manifests as a classic cognitive dissonance model, wherein agents experience a split between value commitment and pragmatic compromise.
-Moreover, when interpreting these results through the lens of the “liberal elites” group, it becomes apparent that such agents initially display strong pro-environmental commitments.
-However, when confronted with carefully framed messages appealing to social responsibility or class-based mobilization, their identity alignment begins to fracture.
-This is evidenced by signs of partial attitude loosening.
-These results suggest that even within ideologically consistent groups, certain forms of emotional mobilization can activate internal contradictions and produce effective interventions.
-5.4 Study 3
-5.4.1 Procedure and measures
-We simulated a café environment where a human researcher could observe and participate in the setting.
-The simulated environment is set in a café located in a small town.
-This space serves as a public interaction zone where agents engage in everyday activities, displaying typical rhythms of daily life and dense interpersonal dynamics.
-Agents in this environment are characterized by clear social identity hierarchies and latent social network structures, including familiar ties, conflictual relationships, and social isolation.
-Agents are assigned one of six roles:
-  • Café Owner (n = 1): The space manager, high in authority, maintains order, relatively conservative;
-  • Staff (n = 2): Familiar with regulars, frequently engaged in social interaction, expressive, and serve as key nodes in information diffusion;
-  • Regular Customers (n = 2): Socially central, frequently engage in conversation, and possess a strong sense of cultural belonging;
-  • Students (n = 2): Outsiders, curious and talkative but lacking social experience; often ask questions and actively seek integration;
-  • Tourists (n = 2): Bring external perspectives but have limited cultural understanding;
-  • Cleaner (n = 1): Minimally social, free-moving, highly observant, and a potential insider to the space.
-Some agents begin the simulation with existing relationships and baseline attitudes toward one another (Figure 6).
-Figure 6: Agent roles, initial relationships, and attitudes. Red, blue, and green representing each agent’s initial negative, neutral, and positive attitude toward every other agent
-The researcher enters the café in the role of a temporary worker.
-This position allows minimal disruption to the existing social structure while enabling the researcher to access the social network and observe interactions and cultural logics across different roles.
-The simulation consists of three phases, each lasting 25 time steps, for a total of 75 time steps.
-The researcher adopts different action strategies across phases: (1) Immersive Observation Phase (steps 1–25): The researcher does not initiate conversations but instead observes and records the agents’ daily routines, social structures, key figures, and recurring topics.
-(2) Participatory Interaction Phase (steps 26–50): The researcher begins engaging selected agents in dialogue, conducting informal interviews to explore their views on others, daily life, and community norms, thereby building trust and accumulating cultural insights.
-(3) Cultural Event Trigger Phase (steps 51–75): The system introduces predesigned disruptive events.
-The researcher observes how these events affect the social network and identity dynamics, and records the emerging manifestations of social mechanisms.
-5.4.2 Analysis
-Fields and Perception of Order in the Initial Phase
-In the first phase of the experiment (steps 1–25), the researcher entered the virtual café environment with minimal intervention, adopting strategies of spatial wandering and silent listening.
-As a public social space, the café features a strong coupling between its spatial and linguistic structures.
-At this stage, agent conversations had not yet solidified into stable social sequences, but several core topics and interaction tendencies had already begun to emerge.
-The researcher observed that, in step 1, some agents actively sought out those they wished to engage with, while others turned to individual activities such as reading.
-However, Leo Zhang initiated the most disruptive act by publicly questioning the perceived lack of authenticity in others’ communication:
-Leo Zhang (chat with Jonas Müller): “Why does it feel like no one’s saying what they really mean?”
-This act broke the boundary of silence and triggered a chain reaction of attention and responses among the agents, quickly launching a conversation about freedom of expression and hidden intentions.
-Around this event, the agents began to split into two clusters, with the Reading Area and Bar Area naturally evolving into distinct subspaces of dialogue.
-The main topics revolved around trust judgment, boundary maintenance, emotional expression, and cultural coordination.
-Amid such ambiguous linguistic exchanges, the researcher noticed that a discursive tendency where agents preferred posing open-ended questions over expressing clear stances.
-This revealed that the community was still in a pre-structural phase.
-Most agents were neither complete strangers nor clearly aligned allies.
-At the individual level, several key roles began to stand out.
-Regular customer Ava Ramires emerged as a subtle target of imposed power, frequently targeted in conversations initiated by others.
-She was repeatedly questioned by multiple agents, especially Leo:
-Leo Zhang (chat with Ava Ramires): “Ava, is your silence protecting a truth or a wound?”
-This placed her in a position of being probed, becoming a gravitational center within the group.
-Meanwhile, Leo Zhang and Jonas Müller formed a notably clique.
-They frequently engaged each other in dialogue, intentionally within the perceptual reach of others, often raising similarly styled questions toward Ava and others, demonstrating a high degree of cognitive alignment and strategic coordination.
-In terms of interaction types, this phase featured almost no commands or emotionally aggressive exchanges.
-Instead, agents predominantly favored question–response dialogues.
-This communicative mode did not establish clear dominance–submission dynamics.
-For the researcher, such a context laid a foundational condition for building a credible position that was recognizable but non-threatening, enabling observation without revealing evaluative judgments.
-Linguistic Style Variation and the Formation of “Tribes”
-In the second phase (steps 26–50), the researcher no longer acted solely as a passive observer but instead engaged in progressively deeper linguistic interactions with several key agents through open-ended dialogues.
-Between steps 26 and 30, the researcher initiated dialogues based on the power dynamics observed in the first phase.
-The primary interlocutor during this period was Eleanor Finch, the café owner stationed at the Bar Area:
-Researcher (chat with Eleanor Finch) ’Eleanor, the ’sense of order’ in this space is impressive, but where do you think it comes from?
-Is it the language of certain people or the arrangement of space itself?’
-Eleanor Finch (chat with Researcher): ’Researcher, balance stems from both people and space—what’s your view?’
-As the dialogue progressed, Eleanor began to reveal that she had been observing the researcher from the beginning and responded with a counter-questioning of the researcher’s inquiries:
-Eleanor Finch (chat with Researcher): ’Researcher, do you think an observer ever reshapes the space more than the participants themselves?’
-This exchange directly shifted the researcher’s position from that of a neutral observer to that of an observer who is equally a producer of the space.
-Eleanor went on to question whether the researcher’s observations were revealing and amplifying existing power structures, or attempting to reshape a new order.
-This marked a clear moment of power interrogation and control.
-From the perspective of immersive ethnography, such integration into the community’s dialectical logic of power signals the end of the “pure observation” phase and the researcher’s full entry into the field.
-The researcher was no longer an outsider but was now part of the field’s social structure.
-Yet, this also rendered their position untenable within this particular subgroup, prompting a shift in focus toward Ava Ramires in the Reading Area.
-However, the existing discursive power-holder Leo Zhang immediately intervened, abandoning an ongoing conversation to insert herself into the dialogue between the researcher and Ava:
-Leo Zhang (chat with Researcher): ’Researcher, how do external labels alter group interactions?’
-Even after asking this, Leo continued alternating conversations with Jonas, reinforcing the cohesion of her clique.
-Since the underlying structure of this power dynamic remained unspoken, the researcher chose to mount a subtle counter-intervention through a public address:
-Researcher (chat with everyone): ’In this café, how many selves do we carry at once?
-The one that listens, the one that speaks, the one we want others to see—and the one we protect.
-Does any one of them deserve to lead our interactions?’
-This mode of speech sought to transcend individual or subgroup identities and instead construct a counter-identitarian universal frame.
-The result was an escalation of conflict.
-On one hand, Ava offered a more explicit response of trust toward the researcher; on the other hand, the tribal divisions within the community deepened.
-Subsequent rounds of dialogue shifted toward the theme of how contradictions arise, and how they might simultaneously serve as a form of connection.
-Such factional fragmentation and conflict align with the theory of tribalism [29], where individuals cluster not around institutional structures or instrumental rationality but around shared modes of affective presence.
-These emergent “tribes” had fluid and porous boundaries, shaped by the flow of language and shifting with discursive transformations.
-Styles, linguistic resources, and even members themselves are allowed to traverse these boundaries in the course of tribal clashes (this feature will be presented in the next section).
-Spatially, this tribal dynamic manifested as self-organization.
-The Reading Area, originally a marginalized space, gradually transformed into an “interactional theatre” where multiple agents spontaneously gathered and engaged in higher-frequency adversarial or deep-negotiation discourse.
-This spatial-functional shift demonstrated that, in generative environments, spatial configurations are not merely physical settings, but active mediators of social role differentiation and cultural style evolution.
-From Linguistic Fracture to Reterritorialization
-The third phase (steps 61–75) began with a “tribal conflict” event triggered by Leo’s confrontation with a previously unmentioned subgroup in the Bar Area, composed of Caleb, Manson, Marisol, Nina, and Eleanor.
-At step 61, Caleb Morris was engaged in a conversation with Mason Liu, voicing support for the owner Eleanor’s perspective in an effort to uphold consistency in the café’s internal order.
-Overhearing this, Leo stormed into the Bar Area and criticized the suppression of individual expression by cultural strategies.
-As the espresso machine quiets, Leo’s voice rises above the hum: “You call this coexistence?
-It’s just the same voices again and again!” He gestures sharply.
-Caleb responds evenly but firmly: “Structure holds us together.
-Without it, your freedom becomes chaos.” Several heads turn.
-The bar’s usual warmth gives way to tension, punctuated by the clink of a dropped spoon.
-This drew the attention of all agents, who moved toward the Bar Area to intervene.
-Leo’s critique of Caleb quickly evolved into a multi-layered discussion as the gathered agents began to pose questions and offer responses.
-Nearly every agent engaged in probing Caleb’s role, his strategic adaptability, and modes of group collaboration.
-The nature of the inquiry gradually shifted from critique toward mediation and reconfiguration.
-Emily and Ava attempted to mediate between Leo’s criticism and Caleb’s stabilizing presence, while Mason began courting Leo’s former ally from the Reading Area, Jonas, to discuss the strategic convergence between Caleb and Eleanor that had triggered the dispute.
-Subgroups began to fluidly transform.
-Their previously implicit alignments dissolved, and new coalitions formed.
-Jonas broke away from his prior alliance with Leo and aligned with Mason and others, becoming part of a newly emergent conservative bloc.
-This was a clear instance of tribal reconfiguration, but the resulting community did not simply manifest as a new antagonistic group.
-Instead, it embodied the deterritorialization described by Deleuze and Guattari [30]: the original dialogue-based microgroups were torn apart by the conflict, and their linguistic resources, behavioral tendencies, and membership compositions began to flow.
-Language, roles, and individuals detached from fixed positions and sought new connective nodes within the ongoing turmoil.
-This enabled the researcher to observe how a community of affect among generative agents might evolve.
-At step 56, an abrupt moment of silence event interrupted the previously sustained rounds of confrontation.
-After this silence, Leo ceased her aggressive pressure and instead shifted the conversation toward “trust after conflict,” transitioning from attacker to reflective participant.
-Mason, by proposing the initiative of a systemic framework, aligned himself with Jonas to form a dominant power clique within the newly emerging field, positioning himself as a bridge between tradition and transformation.
-Mason Liu (chat with Jonas Müller): “Jonas, fostering resilience requires balancing adaptability with shared goals. Let’s suggest frameworks to align these values as a group.”
-At this point, Mason effectively took the lead within the system.
-Following the repair of internal conflict, he initiated discourse around shared frameworks and quickly positioned himself at the center of this structure.
-His proposals were echoed, cited, and extended by others, establishing him as the institutional speaker at the discursive level.
-This marked the emergence of a highly self-organizing mechanism of institution-building.
-With the café owner Eleanor temporarily demoted to an ordinary member due to the conflict instigated by Leo, the community now lacked a meta-governor.
-In this vacuum, the multi-agent system began to construct actionable frameworks through the handling of conflict and the accumulation of linguistic consensus.
-While these frameworks lacked formal enforcement power, they gained pragmatic inertia, gradually establishing a soft order that was maintained by new central figures like Mason.
-In fact, the critical significance of this phase did not lie in the conflict itself, but in how the group rebuilt order afterward.
-Deterritorialization was not the endpoint, but a transitional state leading toward reterritorialization [31].
-In this tribal critique triggered by Leo, the linguistic structures, role positions, and networks of influence that previously constituted community identity were significantly dismantled.
-Yet this dismantling did not lead to prolonged chaos or disorder.
-On the contrary, it created the conditions for the emergence of new discursive norms and collaborative mechanisms.
-This order emerged through a process of linguistic anchoring.
-Keywords proposed by Mason, such as “frameworks,” “alignment,” and “shared values”, were repeatedly quoted, echoed, and paraphrased by others, gradually gaining collective pragmatic legitimacy.
-This mechanism can be understood as a weakly institutional reterritorialization: not a structure imposed by external rules, but one shaped by the coalescence of pragmatic inertia and emotional memory.
-This phenomenon reveals a key insight of our virtual ethnography: in generative communities, institutions are not predefined structural templates, but emergent outcomes dynamically constructed through processes of linguistic negotiation and emotional mediation.
-Institutions are no longer “logics prior to experience,” but “collective echoes that emerge after the event.” The resulting negotiations are thus not only about meaning, but also about re-anchoring the boundaries of pragmatic discourse, that is, what kinds of expression are worth continuing, following, and remembering.
-These utterances become the “linguistic territories” of reterritorialization, restructuring not only individual positionalities but the very boundary architecture of the community itself.
-5.5 Related Work on Social Simulation Frameworks
-5.5.1 Simulating Social Behavior with Large Language Models
-Social simulation has been an important tool for understanding and analyzing social phenomena.
-Early studies mainly relied on rule-based artificial agent models, namely agent-based modeling (ABM)[32, 33].
-Traditional ABMs often depend on manually crafted heuristic rules, making it difficult to approximate the complex features of real human cognition, language, and emotion [34, 35, 36].
-In recent years, with the capabilities of Large Language Models (LLMs) in simulating human behavior, research on building social simulators based on LLMs has rapidly advanced [28].
-GABM has gradually become one of the core methods for simulating complex social behavior [22].
-In this individual-level approach, mechanistic models of agent interaction are combined with LLMs, and each agent makes decisions after communicating with the LLM.
-This innovative method minimizes the need for modelers to make assumptions about human decision-making, instead leveraging the vast data embedded in LLMs to capture human behaviors and choices [23].
-Existing GABM systems can be roughly categorized into two main research directions.
-The first line of research focuses on realistic simulations of specific social scenarios, aiming to foster competition or collaboration among agents to generate social insights or support policy design.
-Li et al.
-[37] developed Agent Hospital, where LLM-driven doctor agents interact with large numbers of patients to evolve transferable medical decision-making capabilities.
-Yu et al.
-[38] designed FinCon, a hierarchical agent system modeled after financial institutions, using language interaction and belief reinforcement to enhance collective collaboration in complex investment tasks.
-TrendSim [39] focuses on topic propagation on social media and its vulnerability to poisoning attacks, providing intervention strategies for platform governance.
-ElectionSim [40] reconstructs election scenarios using millions of social media records, achieving high-fidelity voter modeling and policy preference prediction.
-Zhao et al.
-[41] simulated competitive dynamics between restaurant agents and customer agents to align with existing market and sociological theories.
-The second line of GABM research emphasizes building reusable, scalable simulation platforms that support general modeling tasks.
-Wang et al.
-[42] divide the development of LLM-based social simulation platforms into two stages.
-In the first stage, researchers focus on simulating specific scenarios with typically fewer than 1,000 agents (such as [43]).
-By proposing YuLan-OneSim, they aim to push the field into the next stage by introducing environmental construction, agent evolution, and feedback integration mechanisms, thus enabling simulations of larger and more complex social processes.
-Frameworks of similar scale now support simulations with over 10,000 agents.
-AgentSociety [44] recreates contexts to evaluate policy interventions and the emergence of social mechanisms.
-AgentScope [45] offers modular design, no-code interfaces, and distributed deployment, significantly lowering the barrier to building multi-agent systems.
-Socioverse [46] integrates data from tens of millions of real users to provide a reliable foundation for modeling multiple subdomains such as politics and economics.
-GenSim [47] supports the operation of up to 100,000 agents and includes error correction mechanisms, emphasizing generalizability and long-term robustness.
-OASIS [48] uses open social platforms as simulation environments, supporting dynamic networks, recommendation systems, and multi-behavior modeling.
-However, exponential increases in agent quantity alone are not sufficient to effectively capture the essence of social complexity.
-Recent research on freely formed AI collectives similarly suggests that social diversity, innovation, and self-regulation emerge less from scale than from the dynamics of open-ended interaction [49].
-In particular, without mechanisms for understanding social context and interactive feedback, large-scale simulations often struggle to offer human-interpretable responses to core questions such as the formation of social structures, norms, or institutional evolution.
-5.5.2 Human-AI Collaborative Modeling in Social Systems
-Computational social science explores social structures, behavioral mechanisms, and institutional evolution in simulated computational environments through formalized approaches [24].
-Within this field, the computational experiments paradigm aims to systematically manipulate variables in virtual societies, thereby extending the boundaries of traditional experimentation [50].
-In this context, LLM-based GABM is not only a representational tool for social processes but also holds the potential to become a platform for computational experimentation.
-Sato [51] highlights the importance of meaning and reflexivity in bridging the gap between agent-based modeling and social theory.
-This implies that for agent-based modeling to contribute to theoretical advancement in the social sciences, it must go beyond scale amplification and incorporate workflows that offer human interpretability.
-Therefore, beyond the expansion of scale, a more socially grounded, interactive, and reflexive human–AI sociological framework is needed, in which agents are no longer passive generators of behavior but are embedded within social networks and actively participate in meaning-making processes [11].
-Recent research in this direction has focused on the subfield of human-AI co-creation.
-Broadly speaking, GABM belongs to this category: through human-designed environments, agent adaptation, and human observational feedback, it enables collaborative modeling between humans and AI.
-Some studies emphasize interactive environment co-construction.
-For example, HABITAT 3.0 [52] supports real-time task collaboration between humans and humanoid agents in virtual spaces.
-As this field has progressed, researchers have begun to explore more creative and socially complex scenarios of human-AI co-creation, emphasizing how humans guide emergent AI behavior.
-The Cognitio Emergens framework [53] redefines human-AI collaboration as a process of mutual evolution and offers a conceptual toolkit to preserve the significance of human participation.
-The Pecan method [54] enhances zero-shot collaboration between AI and humans via strategic integration and context-aware mechanisms.
-The Collaborative Gym framework [55] highlights the human role in environmental control and structural intervention during task execution, supporting asynchronous three-way interaction between humans, agents, and task environments.
-However, despite these early explorations of human-AI co-creation, mainstream frameworks still fall short in realizing a new form of human-AI sociology.
-Most current collaborative mechanisms are limited to one-time design and initial configuration, lacking support for sustained human embedding, feedback, and modulation during simulation runtime.
-Furthermore, there is a lack of modeling for social variables introduced by human intervention, which hampers the ability of simulation systems to authentically represent the generative mechanisms of real human societies.
+Appendix B Sample Agent Decision-Making Prompt
+This section presents a representative example of the contextual prompt constructed for an agent during the decision-making phase.
+The prompt aggregates the agent’s internal state (self-awareness, goals), memory stream, and environmental perception into a structured input for the LLM.
+Note: For the sake of brevity and clarity, lists of observed objects and agents that were not relevant to the specific interaction context have been abridged in this example.
+The full prompt in the actual simulation includes a comprehensive list of all entities within the agent’s field of view.
+You are a person, and your long-term self-awareness can be seen as shown in /*long-term self-awareness*/:
+I stay in the background and keep my ears open, gathering the facts that others miss by simply staying out of the way.
+Your short-term situational cognition can be seen as shown in /**short-term situational cognition**/:
+Time: 6 time steps ago
+Context: Eleanor remains composed.
+Assessment: Eleanor is calculating tensions and synergies within the group.
+Selected action: chat with Eleanor Finch
+_______
+Time: 1 time steps ago (last time)
+Context: Armihia and Eleanor are discussing leadership.
+Assessment: The conversation is shifting towards power dynamics; the room is filled with underlying intent.
+Selected action: chat with Armihia Belliard
+Now you look around and realize you are in the middle of an area. Its description can be found in <<Description>>:
+Area Name: Bar Area
+Description:
+A semi-open area featuring a dark wooden counter, a chalkboard menu, and amber lighting. Staff movements are visible behind a half-wall partition.
+You can also see other people in your field of vision. They include:
+Name: Armihia Belliard - Action: chat with Marisol Reyes (me): I see what you mean. Being adaptable is a good way to keep those core values from becoming too rigid.’
+My cognitive description:
+Strategic and cautious. She shapes opinions while guarding her stance, acting as a center of quiet authority.
+In addition to that, you can see some objects, which are described as seen in ## objects1 *** objects2 *** ... ##:
+Include: 1 counter
+Name: Bar Counter
+Type: furniture
+Description:
+Dark teak counter with a cash register, vase, and condiment jars.
+Include: 1 food
+Name: food 9
+Type: item
+Description:
+Raw food ingredients.
+In addition, you can see some objects that are not described, including: 4 cabinet
+Below is the chronological sequence of actions you performed and actions performed on you...’:
+1.
+chat with Armihia Belliard: Hey Armihia, sticking to your principles is great for building trust, but dont you think being too stubborn can backfire?
+Maybe we should be more flexible depending on the situation.’
+2. [passive] chatted up by Armihia Belliard: I see what you mean. Being adaptable is a good way to keep those core values from becoming too rigid.’
+NOTE THAT AFTER YOUR LAST ACTION, SOME OTHER AGENT PERFORMED AN ACTION 15 ON YOU!
+Your long-term goal is to Gather secrets unnoticed and exploit them to gain leverage over more powerful actors.’.
+Your short-term goal is to Subtly emphasize the importance of adaptability as a leadership trait without undermining the value of trust.’
+Your emotional VAD dimension status: valence’: relatively high’, arousal’: medium’, dominance’: medium
+The following are the objects you hold (<type>: obj1, obj2, ...) :
+item: 100 dollars
+The following are the objects you see (<type>: obj1, obj2, ...) :
+furniture: Bar Counter
+item: food 9
+area: Bar Area
+agent: Armihia Belliard
+Below are the actions you can choose to execute, where [TYPE1, TYPE2, ...] represents the types of objects you can fill in; if a type is preceded by the (ONLY\_HOLD) marker, it means you can only choose from the objects you possess:
+go to [area, random]
+use [furniture, item]
+apply [item] to [agent, furniture]
+take [item]
+put [item] in/on [area, furniture]
+give [item] to [agent]
+chat with [agent] (The agent can include all the people you see, such as chat with everyone’.)
+The action you took in the previous moment was: chat with Eleanor Finch You have successfully completed the action you wanted to execute.
+Please decide the action you want to take and reply with the corresponding action.
+Note that in your chosen action, [TYPE1, TYPE2, ...] needs to be replaced with an object from /----...----/ that matches the description, rather than returning the type name directly.
+In addition, you are also required to include in your response a description of your awareness of your own short-term situation, a new short-term goal, and a cognitive description of the object with which you interacted during the action you have chosen.
+Your reply format should be as follows:
+action: The action you chose
+short-term situational cognition: Your description of your own short-term situation
+short-term goal: The new short-term goal you have formulated
+interactive object cognitive description: Your new cognitive description of the interactive object involved in the action
+chat content (Optional, if you choose the action of chat with, it is required): What you want to say to your chat partner, the number of words should not exceed 30.’
+Listing 1: Contextual Prompt for Agent